### v1

# On Nontrivial Winning and Losing Parameters of Schmidt Games

Abstract.
In this paper we completely describe the winning and losing conditions different from the only “trivial” conditions known before. In other words, we solve the open question of finding a complete nontrivial Schmidt diagram. In addition, we give the new bounds for two family of sets: one related to frequencies of digits in base-$2$ expansions, and one connected to the set of the badly approximable numbers.
1. Set-up and Introduction: Schmidt Games
Wolfgang Schmidt first introduced Schmidt games in [S66] in order to study the set of badly approximable numbers. Winning sets in these games have special properties, such as a full Hausdorff dimension; as such, Schmidt games have been used to prove results in number theory and dynamical systems. Here, we briefly describe the setup of a Schmidt $(\alpha,\beta)$-game. Throughout this paper we work with the Schmidt game on $\mathbb{R}$, so for simplicity we give the definitions only in this special case.
Denote by $I$ the open unit square:
$I\overset{\rm{def}}{=}\{(\alpha,\beta):0<\alpha,\beta<1\}=(0,1)\times(0,1)\,.$
Then, fix a set $S\subset{\mathbb{R}}$ and $(\alpha,\beta)\in I$. Two players, Alice and Bob, can play a Schmidt game on $S$, taking alternating turns, as follows: Bob begins by choosing an interval $B_{0}$. Given an interval $B_{i}$, Alice chooses an interval $A_{i}\subset B_{i}$ such that $|A_{i}|=\alpha|B_{i}|$. Similarly, given an interval $A_{i}$, Bob chooses an interval $B_{i+1}\subset A_{i}$ such that $|B_{i+1}|=\beta|A_{i}|$. Let
$x_{\infty}\overset{\rm{def}}{=}\bigcap_{k=1}^{\infty}B_{k}=\bigcap_{k=1}^{\infty}A_{k}.$
We call $S$ $(\alpha,\beta)$-winning if, regardless of how Bob chooses his intervals, Alice can play in such a way that $x_{\infty}\in S$. For every $S$, define the Schmidt diagram $D(S)$ of $S$ as the set of all pairs $(\alpha,\beta)\in I$ such that $S$ is $(\alpha,\beta)$-winning. In his original paper, Schmidt proved that there are two “trivial zones” of Schmidt diagrams:
Lemma 1.1.
[S66]
  • (a)
    If $S$ is dense, then $D(S)$ contains the set
    $\check{D}:=\big{\{}(\alpha,\beta)\in I:\beta\geq(2-\alpha)^{-1}\big{\}},$
    with equality if $S$ is countable (see Figure 2).
  • (b)
    If $S\subsetneq{\mathbb{R}}$, then $D(S)$ is contained in the set
    $\hat{D}:=\big{\{}(\alpha,\beta)\in I:\beta>2-\alpha^{-1}\big{\}},$
    with equality if the complement of $S$ is countable (see Figure 2).
Figure 1. $\check{D}$ (blue)
Figure 1. $\check{D}$ (blue)
Figure 2. $\hat{D}$ (blue)
There are only four sets which were known to coincide with $D(S)$ for some $S$: $\emptyset$, $\check{D}$, $\hat{D}$, and $I$. Our goal with this project was to find some set $S$ such that $\check{D}\subsetneq D(S)\subsetneq\hat{D}$ and to completely describe $D(S)$. It is worth noting that this problem was investigated in an earlier PRIMES paper, which investigated sets relating to particular decimal expansions [KZT12]. In the aforementioned paper, some nontrivial statements about Schmidt diagrams were obtained, but no $D(S)$ different from the four diagrams discussed above was completely described.
2. Main Results
One of the main results of this paper is a complete description of a Schmidt Diagram different from $\emptyset$, $\check{D}$, $\hat{D}$ and $I$. More specifically, we prove the following theorem.
Theorem 2.1.
There exists a set $\mathscr{S}$ such that $D(\mathscr{S})=\{(\alpha,\beta):0<\alpha<\beta<1\}$ (see Figure 3).
Figure 3. $D(\mathscr{S})$ (blue)
We develop the techniques for this in Section 3, and the more general statements about sets with such a diagram are given in Theorem 3.10 and Corollary 3.11. We construct a special family of examples in Section 4.
In Section 5, we describe some winning and losing conditions for families of sets connected with digit frequencies in the dyadic expansion of real numbers. We prove a general case in Theorem 5.3 and a special symmetric case in Corollary 5.6. In Section 6 we give some new results about winning and losing conditions for families of sets connected with badly approximable numbers in base $B$ expansion: specifically in Theorems 6.4, 6.5, and 6.6.
The bounds for the winning and losing zones we describe in Sections 5 and 6 are not optimal; however, they give some estimates and shed some light on how their respective Schmidt diagrams can look like. Figures 5 and 5 show two examples of the possible bounds (where the blue zones are subsets of the corresponding Schmidt diagram and the red zones are subsets of its complement). The sets $D_{c}^{-}$ used in the first example will be defined later in Section 5 and the sets $\operatorname{\widetilde{2-BA}}(c)$ will be defined in Section 6.
Figure 4. Bounds for $D(D_{c}^{-})$ with $c=0.7$
Figure 4. Bounds for $D(D_{c}^{-})$ with $c=0.7$
Figure 5. Bounds for $D(\operatorname{\widetilde{2-BA}}(c))$ with $c=0.035$
3. General strategies in the Schmidt’s game
This section gives some general statements concerning strategies in Schmidt’s game.
Let us start with a classical result due to Schmidt:
Lemma 3.1.
[S66] If a set $S$ is $(\alpha,\beta)$-winning, $\alpha\beta=\alpha^{\prime}\beta^{\prime}$, and $\alpha>\alpha^{\prime}$, then $S$ is also $(\alpha^{\prime},\beta^{\prime})$-winning.
Sometimes for simplicity we will also refer to existence the of positional strategies, i.e. to the following classical statement:
Theorem 3.2.
[S66] If a set $S$ is winning, there exists a positional winning strategy; that is, a strategy that only depends on the most recent interval played and not the prior history of moves.
The next four statements are the technical steps in proving Lemma 3.7; however, they are rather general and can be used separately in the further research.
Lemma 3.3.
Fix $m\in\mathbb{N}$. Let $D\subseteq\mathbb{R}$ be some set with a property that $m^{k}D=m^{k}+D=D$ for any integer $k$. Suppose Bob has a strategy for $(\alpha,\beta)$-game with the set $D$ that allows him to win, and the first ball in this strategy is $B_{0}$. Then:
  1. (1)
    For any $k\in\mathbb{Z}$ Bob has a winning strategy for $(\alpha,\beta)$-game with the set $D$ starting with the ball $m^{k}B_{0}$;
  2. (2)
    For any $k\in\mathbb{Z}$ Bob has a winning strategy for $(\alpha,\beta)$-game with the set $D$ starting with the ball $m^{k}+B_{0}$.
(Here $m^{k}B_{0}=\{m^{k}x:\,\,\,x\in B_{0}\}$ and $m^{k}+B_{0}=\{m^{k}+x:\,\,\,x\in B_{0}\}$).
Proof.
  1. (1)
    Suppose Bob starts the game with $B_{0}^{\prime}=m^{k}B_{0}$, and Alice plays some ball $A_{0}^{\prime}\subset B_{0}^{\prime}$. Let us define $A_{0}=m^{-k}A_{0}^{\prime}\subseteq B_{0}$. Now let $B_{1}\subset A_{0}$ be the ball that Bob would play after $A_{0}$ according to his winning strategy. Let Bob play $B_{1}^{\prime}=m^{k}B_{1}\subset A_{0}^{\prime}$.
    After $n$ steps, analogously, suppose Alice plays some ball $A_{n}^{\prime}\subset B_{n}^{\prime}$. Define $A_{n}=m^{-k}A_{n}^{\prime}$, and define $B_{n+1}$ as the ball Bob would play after $A_{n}$; then let Bob play $B_{n+1}^{\prime}=m^{k}B_{n+1}\subset A_{n}^{\prime}$.
    It is easy to see that $\bigcap\limits_{i=0}^{\infty}B_{i}^{\prime}=m^{k}\bigcap\limits_{i=0}^{\infty}B_{i}=m^{k}x$, where $x$ is the outcome of some game with the starting ball $B_{0}$, i.e. $x\notin D$, therefore $m^{k}x\notin D$ and Bob wins.
  2. (2)
    The strategy and the argument is analogous; we just take $A_{i}=A_{i}^{\prime}-m^{k}$ and $B_{i}^{\prime}=B_{i}+m^{k}$ and use the fact that $x\in D$ if and only if $m^{k}+x\in D$.
Lemma 3.4.
Let $S\subseteq\mathbb{R}$ be any set. Suppose Bob has a strategy for $(\alpha,\beta)$-game with the set $S$ that allows him to win, and the first ball in this strategy is $B_{0}$. Then for any $(\alpha^{\prime},\beta^{\prime})$ such that $\beta^{\prime}<\beta$ and $\alpha^{\prime}\beta^{\prime}=\alpha\beta$ Bob has a winning strategy starting from $B_{0}$ in $(\alpha^{\prime},\beta^{\prime})$-game.
Proof.
Suppose Bob starts $(\alpha^{\prime},\beta^{\prime})$-game with the ball $B_{0}^{\prime}=B_{0}$, and Alice plays some ball $A_{0}^{\prime}$. Let $A_{0}$ be the ball of size $|A_{0}|=\alpha|B_{0}|=\frac{\alpha}{\alpha^{\prime}}|A_{0}^{\prime}|<|A_{0}^{\prime}|$ centered at the center of $A_{0}^{\prime}$; suppose $B_{1}$ is the ball Bob would play after $A_{0}$ in $(\alpha,\beta)$-game. Let Bob play the ball $B_{1}^{\prime}=B_{1}$ after $A_{0}^{\prime}$ in $(\alpha^{\prime},\beta^{\prime})$-game; indeed, $B_{1}^{\prime}\subset A_{0}\subseteq A_{0}^{\prime}$ and $|B_{1}^{\prime}|=\beta|A_{0}|=\alpha\beta|B_{0}^{\prime}|$, so Bob can make such a move.
After $n$ steps, analogously, suppose Alice plays some ball $A_{n}^{\prime}\subset B_{n}^{\prime}$. Define $A_{n}$ to be the ball of size $|A_{n}|=\alpha|B_{n}|=\frac{\alpha}{\alpha^{\prime}}|A_{n}^{\prime}|<|A_{n}^{\prime}|$ centered at the center of $A_{n}^{\prime}$, and define $B_{n+1}$ as the ball Bob would play after $A_{n}$; then let Bob play $B_{n+1}^{\prime}=B_{n+1}\subset A_{n}^{\prime}$. One has $\bigcap\limits_{i=0}^{\infty}B_{i}^{\prime}=\bigcap\limits_{i=0}^{\infty}B_{i}=x$ where $x$ is the outcome of some game with the starting ball $B_{0}$, i.e. $x\notin S$, therefore Bob wins.
Lemma 3.5.
Fix $m\in\mathbb{N}$. Let $D\subseteq\mathbb{R}$ be some set with a property that $m^{k}D=m^{k}+D=D$ for any integer $k$. Let parameters $(\alpha,\beta)$ be such that $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, and suppose Bob wins $(\alpha,\beta)$-game with $D$. Then for any $L>0$ and any $\varepsilon>0$ there exists a winning strategy in this game for Bob starting with $B_{0}$ with
$L\leq|B_{0}|<(1+\varepsilon)L.$
Proof.
Suppose the ball Bob starts his winning strategy with has size $l$. We denote $\xi=-\log_{\alpha\beta}m$ and $\chi=\log_{\alpha\beta}(\frac{L}{l})$, and let $\varepsilon^{\prime}=\log_{\alpha\beta}(1+\varepsilon)$.
As $\xi$ is irrational, the sequence $k\xi+\chi+\varepsilon^{\prime}$ is uniformly distributed modulo 1, so there exists $k$ such that $0<\{k\xi+\chi+\varepsilon^{\prime}\}\leq\varepsilon^{\prime}$. Let $n=\lfloor k\xi+\chi+\varepsilon^{\prime}\rfloor$; in this case
$\log_{\alpha\beta}L+k\xi\leq n+\log_{\alpha\beta}l<\log_{\alpha\beta}L+k\xi+\varepsilon^{\prime}$
and
$\frac{L}{m^{k}}\leq(\alpha\beta)^{n}l=|B_{n}|<(1+\varepsilon)\frac{L}{m^{k}},$
where $B_{n}$ is any ball Bob can play using his strategy. Fix some possible $B_{n}$; if it occurs in the game, Bob can start the game with this $B_{n}$. By Lemma 3.3 Bob can start also with $B_{0}=m^{k}B_{n}$ and win. It remains to notice that $|B_{0}|=m^{k}|B_{n}|$ and
$L\leq|B_{0}|<(1+\varepsilon)L.$
Lemma 3.6.
Let $S\subseteq\mathbb{R}$ be any set. Suppose Bob has a strategy for $(\alpha,\beta)$-game with the set $S$ that allows him to win, and the first ball in this strategy is $B_{0}$. Then for any $\varepsilon>0$ and for any $B_{0}^{\prime}\subseteq B_{0}$ with $|B_{0}^{\prime}|\geq\frac{|B_{0}|}{1+\varepsilon}$, Bob has a winning strategy starting from $B_{0}^{\prime}$ in $((1+\varepsilon)\alpha,\frac{\beta}{1+\varepsilon})$-game.
Proof.
Fix $\varepsilon>0$. Suppose Bob starts $(\alpha^{\prime},\beta^{\prime})$-game (where $\alpha^{\prime}=(1+\varepsilon)\alpha$ and $\beta^{\prime}=\frac{\beta}{1+\varepsilon}$) with some $B_{0}^{\prime}$ such that $|B_{0}^{\prime}|=\frac{|B_{0}|}{1+\varepsilon}$.
Suppose Alice chooses some $A_{0}^{\prime}\subseteq B_{0}^{\prime}\subseteq B_{0}$; one has $|A_{0}^{\prime}|=\alpha^{\prime}|B_{0}^{\prime}|=(1+\varepsilon)\alpha\frac{|B_{0}|}{1+\varepsilon}=\alpha|B_{0}|$. Let $A_{0}=A_{0}^{\prime}$, and let $B_{1}$ be the ball Bob would play after $A_{0}$ in $(\alpha,\beta)$-game. Now, define $B_{1}^{\prime}$ as a ball centered at the center of $B_{1}$ having $|B_{1}^{\prime}|=\frac{|B_{1}|}{1+\varepsilon}$, and let Bob play $B_{1}^{\prime}$ after $A_{0}^{\prime}$ in our $(\alpha^{\prime},\beta^{\prime})$-game.
Analogously, let $A_{n}=A_{n}^{\prime}$, $B_{n+1}$ the ball Bob would play after $A_{n}$ in $(\alpha,\beta)$-game and $B_{n+1}^{\prime}$ a ball centered at the center of $B_{n+1}$ having $|B_{n+1}^{\prime}|=\frac{|B_{n+1}|}{1+\varepsilon}$; Bob can play $B_{n+1}^{\prime}$ after any possible $A_{n}^{\prime}$.
One can see that $\bigcap\limits_{i=0}^{\infty}B_{i}^{\prime}=\bigcap\limits_{i=0}^{\infty}B_{i}=x$ where $x\notin S$, as it is the outcome of an $(\alpha,\beta)$-game where Bob followed his winning strategy. Therefore Bob wins.
Finally, for an arbitrary $|B_{0}^{\prime}|\geq\frac{|B_{0}|}{1+\varepsilon}$ let us note that $|B_{0}^{\prime}|=\frac{|B_{0}|}{1+\varepsilon^{\prime}}$ for $\varepsilon^{\prime}\leq\varepsilon$, which means that Bob can win $((1+\varepsilon^{\prime})\alpha,\frac{\beta}{1+\varepsilon^{\prime}})$-game on $S$ starting from $B_{0}^{\prime}$. Lemma 3.4 completes the proof. ∎
Lemma 3.7.
Fix $m\in\mathbb{N}$. Let $D\subseteq\mathbb{R}$ be some set with a property that $m^{k}D=m^{k}+D=D$ for any integer $k$. Let parameters $(\alpha,\beta)$ be such that $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, and suppose Bob wins $(\alpha,\beta)$-game with $D$. Then for any $\beta^{\prime}<\beta,\,\alpha^{\prime}=\frac{\alpha\beta}{\beta^{\prime}}$, and any closed ball $T$ Bob can win $(\alpha^{\prime},\beta^{\prime})$-game starting with any ball $T$ as his initial ball.
Proof.
Fix an arbitrary ball $T$. Also, fix $\varepsilon>0$, and let $L=|T|$. By Lemma 3.5 Bob has a winning strategy in $(\alpha,\beta)$-game with a starting ball $B_{0}$ s.t. $(1+\frac{\varepsilon}{2})L\leq|B_{0}|<(1+\varepsilon)L$. Take $\varepsilon^{\prime}<\varepsilon$ such that $|B_{0}|=(1+\varepsilon^{\prime})L$.
Now, let $k$ be such that $\frac{1}{m^{k}}<\frac{1}{2}\varepsilon^{\prime}L=\frac{1}{2}(|B_{0}|-|T|)$; let $b_{0}$ be the center of $B_{0}$ and $t$ the center of $T$. There exists $q\in\mathbb{Z}$ such that $|\frac{q}{m^{k}}+b_{0}-t|<\frac{1}{m^{k}}$. Consider the ball $B_{0}^{\prime}=B_{0}+\frac{q}{m^{k}}$: by Lemma 3.3, Bob has a winning strategy in $(\alpha,\beta)$-game with a starting ball $B_{0}^{\prime}$. In addition, note that $T\subseteq B_{0}^{\prime}$ and $|T|=\frac{|B_{0}^{\prime}|}{1+\varepsilon^{\prime}}$. By Lemma 3.6, Bob has a winning strategy in $((1+\varepsilon)\alpha,\frac{\beta}{1+\varepsilon})$-game with a starting ball $T$.
Now, for $(\alpha^{\prime},\beta^{\prime})$ as in Lemma it remains to take $\varepsilon$ under the condition $(1+\varepsilon)\alpha=\alpha^{\prime}$. ∎
The discussion above implies the following statement:
Lemma 3.8.
Let $m\in\mathbb{N}$ and $D,Q\subseteq\mathbb{R}$ be some sets with the following properties:
  • $m^{k}D=m^{k}+D=D$ for any integer $k$.
  • $D\sqcup(-D)\sqcup Q=\mathbb{R}$.
Then:
  1. (1)
    If $\alpha\geq\beta$, then the set $D$ is $(\alpha,\beta)$-losing.
  2. (2)
    If $\beta>\alpha$ and $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, then the set $D\sqcup Q$ is $(\alpha,\beta)$-winning.
Proof.
  1. (1)
    Suppose $\alpha\geq\beta$ and $D$ is $(\alpha,\beta)$-winning. In this case Alice has a positional winning strategy described by the function $A(I)$ (where $I\subset\mathbb{R}$ is a closed interval) for this game. Consider $(\beta,\alpha)$-game on the set $D$ and let Bob play the following strategy:
    • Bob chooses $B_{0}=A([0,1])$.
    • If Alice plays the interval $A_{n}$ at the $n$-th step, Bob plays $B_{n+1}=-A(-A_{n})$.
    Suppose $x=\bigcap\limits_{i=0}^{\infty}B_{i}$ is the outcome of the game. Then $-x=\bigcap\limits_{i=0}^{\infty}A(-A_{n})$ is the outcome of the $(\alpha,\beta)$-game where Bob plays the balls $[-1;0];-A_{0},-A_{1},\ldots$ and Alice plays using her winning strategy $A(I)$. Therefore, $-x\in D$, which means $-x\notin D$ and Bob wins.
    It means that the point $(\alpha,\beta)$ is winning and the point $(\beta,\alpha)$ is losing for the set $D$, which contradicts Lemma 3.1. Thus, the point $(\alpha,\beta)$ is losing.
  2. (2)
    The idea of the proof is the same as in (a). Suppose $D\sqcup Q$ is $(\alpha,\beta)$-losing; Take $\varepsilon>0$ under the condition $(1+\varepsilon)^{2}<\frac{\beta}{\alpha}$. Then by Lemma 3.7 for any closed ball $T$ Bob has a winning strategy for $\left(\alpha(1+\varepsilon),\frac{\beta}{(1+\varepsilon)}\right)$-game starting from $T$. We use this statement to construct a $\left(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon)\right)$-winning strategy for Alice.
    Suppose Bob starts some $(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon))$-game with the ball $B_{0}$. Then Alice picks some ball $A_{0}\subseteq B_{0}$ without any restrictions. If Bob plays the ball $B_{n}$ on $n$-th step, Alice responds by playing $A_{n}$ where $-A_{n}$ is the ball Bob would play in $(\alpha(1+\varepsilon),\frac{\beta}{(1+\varepsilon)})$-game which started from $-A_{0}$ after the balls $-A_{0},-B_{1},\ldots,-B_{n}$ (in this game $-A_{i}$ are Bob’s balls and $-B_{i}$ are Alice’s). Let $x$ be the outcome of this $(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon))$-game; then $-x$ is the outcome of $(\alpha(1+\varepsilon),\frac{\beta}{(1+\varepsilon)})$-game Bob played with his winning strategy, i.e. $-x\in(-D)$ and $x\in D\subseteq D\sqcup Q$. It shows that $(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon))$ is a winning point for $D\sqcup Q$.
    But one has $\alpha<\alpha(1+\varepsilon))<\frac{\beta}{(1+\varepsilon)}$, which by Lemma 3.1 contradicts our initial assumption. Therefore $D\sqcup Q$ is $(\alpha,\beta)$-winning.
Now, the goal of the next part is to find the condition under which we can remove dyadic rationals from the set $D$ so that it remains winning for the parameters above.
We start with the following almost obvious statement:
Lemma 3.9.
Suppose $S$ is an $(\alpha,\beta)$-winning set and for some $\xi\in\mathbb{R}$ one has $S+\xi\subseteq D$. Then the set $D$ is $(\alpha,\beta)$-winning.
Proof.
If $A(I)$ was Alice’s winning positional strategy on the set $S$, then $A(I-\xi)+\xi$ is a winning positional strategy on the set $D$. ∎
Now, we use the lemmas we have to prove the following statement:
Theorem 3.10.
Fix $m\in\mathbb{Z}$. Let $D\subseteq\mathbb{R}$ be some set with the following properties:
  • $m^{k}D=m^{k}+D=D$ for any integer $k$.
  • $D\sqcup(-D)\sqcup Q=\mathbb{R}$, where $Q\subseteq\mathbb{Q}$ is some set.
  • $D+\xi\subseteq D$ for some $\xi\notin\mathbb{Q}$.
Then:
  1. (1)
    If $\alpha\geq\beta$, then the set $D$ is $(\alpha,\beta)$-losing.
  2. (2)
    If $\beta>\alpha$ and $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, then the set $D$ is $(\alpha,\beta)$-winning.
Proof.
  1. (1)
    follows from Lemma 3.8.
  2. (2)
    By Lemma 3.8 we know that the set $D\sqcup Q$ is $(\alpha,\beta)$-winning. As $Q\subseteq\mathbb{Q}$, one has $D\sqcup Q+\xi\subseteq D$ for any $\xi\notin\mathbb{Q}$. By Lemma 3.9 then, $D$ is $(\alpha,\beta)$-winning for $\beta>\alpha$.
Corollary 3.11.
Let $D\subseteq\mathbb{R}$ be some set with the following properties:
  • $m_{i}^{k}D=m_{i}^{k}+D=D$ for any integer $k$ and some $m_{1},\,m_{2}\in\mathbb{N}$ with property $\log_{m_{1}}(m_{2})\notin\mathbb{Q}$;
  • $D\sqcup(-D)\sqcup Q=\mathbb{R}$, where $Q\subseteq\mathbb{Q}$ is some set;
  • $D+\xi\subseteq D$ for some $\xi\notin\mathbb{Q}$.
If $\beta>\alpha$, then the set $D$ is $(\alpha,\beta)$-winning.
Indeed: for any parameters $(\alpha,\beta)$ in this case either $\log_{m_{1}}(\alpha\beta)$ or $\log_{m_{2}}(\alpha\beta)$ is irrational, so by Theorem 3.10 $D$ is $(\alpha,\beta)$-winning.
4. Example: sets with a nontrivial diagram
Theorem 3.10 and Corollary 3.11 give the complete description of the nontrivial Schmidt diagram for sets with special properties. The goal of this section is to give an example of a set satisfying these properties, thus proving Theorem 2.1.
Fix $\xi\notin\mathbb{Q}$. Then, we define an auxiliary set $\Xi$ by $\Xi=\{r+q\xi:\,\,r,q\in\mathbb{Q},\,q>0\}$. One can see that $-\Xi=\{r+q\xi:\,\,r,q\in\mathbb{Q},\,q<0\}$ is disjoint with $\Xi$, and one has $\Xi+\xi\subseteq\Xi$ by construction.
Now, let us focus on the set $E=\mathbb{R}\setminus\Big{(}\mathbb{Q}\cup\Xi\cup(-\Xi)\Big{)}$. We will divide it into the union of the sets $\overline{x}=\{r+q\xi+px:\,\,r,q,p\in\mathbb{Q},\,\,p\neq 0\}$ over $x\in E$. Easy to see that $\overline{x}$ is the equivalence class under relation $\sim$ defined as follows: $x\sim y$ if $y=r+q\xi+px$ with the restrictions as above, therefore $E$ is the disjoint union of the sets $\overline{x}$ over pairwise non-equivalent $x$.
Next, let $\underline{x}=\{r+q\xi+px:\,\,r,q,p\in\mathbb{Q},\,\,p>0\}$. Then $\overline{x}=\underline{x}\sqcup\underline{-x}$. Indeed: if $\underline{x}\,\text{and}\,\underline{-x}$ intersect, then $x\in\Xi\sqcup(-\Xi)$, which we do not consider. Finally, note that $\underline{-x}=-\underline{x}$.
Let us enumerate all the sets $\overline{x}$ by some index set $\mathscr{A}$ and denote the corresponding set $E_{\alpha}$ where $\alpha\in\mathscr{A}$. Choose some $x_{\alpha}\in A_{\alpha}$ and denote $\underline{x_{\alpha}}=C_{\alpha}$; then $C_{\alpha}\sqcup(-C_{\alpha})=E_{\alpha}$. Therefore for $C=\bigsqcup\limits_{\alpha\in\mathscr{A}}C_{\alpha}$ one has $C\sqcup(-C)=E$; by construction also $C+\xi_{j}=C$ (as this works for each of $\underline{x}$).
Finally, we define $\mathscr{S}=\mathscr{S}(\xi)=C\sqcup\Xi$. Then $-\mathscr{S}=(-C)\sqcup(-\Xi)$, and one has $\mathscr{S}\sqcup(-\mathscr{S})\sqcup\mathbb{Q}=\mathbb{R}$. One also has:
  • $m^{k}\mathscr{S}=m^{k}+\mathscr{S}=\mathscr{S}$ for any integer $k$ and any $m\in\mathbb{N}$;
  • $\mathscr{S}+\xi\subseteq\mathscr{S}$.
Therefore, by Theorem 3.10 and Corollary 3.11, $\mathscr{S}$ is $(\alpha,\beta)$-winning for $\beta>\alpha$ and $(\alpha,\beta)$-losing for $\beta\leq\alpha$. This proves Theorem 2.1.
5. Digit frequencies
Next, we will investigate a separate but similar construction regarding frequencies of digits in base-$2$ expansions of real numbers. It is known that these sets are closely related to density and dimension, and hence they have been studied in the past (see [V59, FFNS21]).
First, some setup. Consider the base-$2$ expansions of the form $x=x_{0}.x_{1}x_{2}\cdots$ where $x_{0}$ is an integer and $x_{i}\in\{0,1\}$ are the digits in the base-$2$ expansion of $x$. In this paper, if $x$ can be represented by two seperate expansions, we take the lexicographically minimal expansion (for example, $0.100\ldots$ would instead be written as $0.011\ldots$). We define
$d^{+}(x,j)=\limsup_{k\rightarrow\infty}\frac{\#\{1\leq i\leq k:x_{i}=j\}}{k}$
and similarly,
$d^{-}(x,j)=\liminf_{k\rightarrow\infty}\frac{\#\{1\leq i\leq k:x_{i}=j\}}{k}.$
Define the sets
$D^{+}_{c}$ :=$:=\{x\in{\mathbb{R}}:d^{+}(x,0)>c\},$
$D^{-}_{c}$ :=$:=\{x\in{\mathbb{R}}:d^{-}(x,0)>c\}.$
The following is known about the Schmidt diagrams of these sets.
Theorem 5.1.
[FFNS21] The set $D^{+}_{c}$ is winning for $\log(8\alpha)/\log(\alpha\beta)>c$ and losing for $\log(\alpha/8)/\log(\alpha\beta)\leq c$.
The results of our research expand on that bound. To do so, we require the following lemma, known as the Equidistribution Theorem:
Lemma 5.2.
[B09] If $\alpha$ is irrational, then the sequence $\alpha,2\alpha,3\alpha,\ldots$ is uniformly distributed modulo $1$.
Theorem 5.3.
Suppose $\log_{2}{\alpha\beta}\notin\mathbb{Q}$. Then, the sets $D^{\pm}_{c}$ are winning for $\alpha\leq\frac{1}{4}$ and $\log(4\alpha)/\log(\alpha\beta)>c$, and losing for $\beta\leq\frac{1}{4}$ and $\log(\alpha/4)/\log(\alpha\beta)\leq c$.
In the special case the bounds given by Theorem 5.3 are presented on Figure 5.
Proof.
We will show the winning strategy for Alice.
Suppose Bob begins by playing an interval with radius $r_{0}$. Alice will play in the following way: suppose Bob plays some interval $B_{t}$. Define $k_{t}=-\left\lfloor\log_{2}{(\alpha\beta)^{t}r_{0}}\right\rfloor$, and pick the smallest $m\in\mathbb{Z}$ such that $m/2^{k_{t}}\in B_{t}$. Then, it follows that the dyadic interval $\left(\frac{m}{2^{k_{t}}},\frac{m+1}{2^{k_{t}}}\right)\subset B_{t}$. Then, Alice plays $A_{t}=\left(\frac{m}{2^{k_{t}}},\frac{m}{2^{k_{t}}}+2\alpha(\alpha\beta)^{t}r_{0}\right)$. Define $l_{t}$ such that $k_{t}+l_{t}=-\left\lfloor\log_{2}{\alpha(\alpha\beta)^{t}r_{0}}\right\rfloor-2$. Then, Alice ensures that for any $x\in A_{t}$, the $(k_{t}+1)$th digit to the $(k_{t}+l_{t})$th digit is equal to 0. In other words, she guarantees $l_{t}$ zeroes on this turn. For convenience, let $a=\log_{2}\alpha$, $b=\log_{2}\beta$, and $\rho=\log_{2}r_{0}$; note that $a+b\notin\mathbb{Q}$. We have
$l_{t}$ $=(k_{t}+l_{t})-k_{t}$
$=\left\lfloor(a+b)t+\rho\right\rfloor-\left\lfloor a+(a+b)t+\rho\right\rfloor-2$
$=(a+b)t+\rho-(a+(a+b)t+\rho)-\{(a+b)t+\rho\}+\{a+(a+b)t+\rho\}-2$
$=-(a+2)+\{a+(a+b)t+\rho\}-\{(a+b)t+\rho\}.$
This implies that by the Equidistribution Theorem,
$\lim\limits_{n\rightarrow\infty}\frac{\sum_{i=0}^{n-1}l_{i}}{n}$ $=-(a+2).$
Now, notice that for $k_{n}\leq k<k_{n+1}$ one has
#$\frac{\#\{1\leq i\leq k:x_{i}=0\}}{k}=\frac{\#\{1\leq i\leq k_{n}:x_{i}=0\}}{k_{n}}\cdot\frac{k_{n}}{k}+\frac{\#\{k_{n}+1\leq i\leq k:x_{i}=0\}}{k},$
which implies
#$\frac{\#\{1\leq i\leq k_{n}:x_{i}=0\}}{k_{n}}\cdot\frac{k_{n}}{k_{n+1}}\leq\frac{\#\{1\leq i\leq k:x_{i}=0\}}{k}.$
It is easy to see that $k_{n+1}-k_{n}$ is bounded, therefore $\lim\limits_{n\rightarrow\infty}\frac{k_{n}}{k_{n+1}}=1$, and
$d^{-}(x,0)$ $=\liminf\limits_{k\rightarrow\infty}\frac{\#\{1\leq i\leq k:x_{i}=0\}}{k}$
$\geq\liminf\limits_{n\rightarrow\infty}\frac{\#\{1\leq i\leq k_{n}:x_{i}=0\}}{k_{n}}\cdot\frac{k_{n}}{k_{n+1}}$
$\geq\lim\limits_{n\rightarrow\infty}\frac{\sum_{i=0}^{n-1}(l_{i})}{1-\log_{2}((\alpha\beta)^{n}r_{0})}\cdot\frac{k_{n}}{k_{n+1}}$
$=\frac{-(a+2)}{-(a+b)}$
$=\frac{\log(4\alpha)}{\log(\alpha\beta)}>c$
as desired. Therefore, Alice wins. The losing condition is analogous, so we omit it.
For the proofs below, we will use the following
Lemma 5.4.
Let $S\subseteq\mathbb{N}$ be an infinite set of density zero. Let $d^{\pm}(x,0)=c$; then
$d^{\pm}\left(x+\sum\limits_{j\in S}\frac{1}{2^{j}},0\right)\geq c.$
Proof.
For $x\in\mathbb{Z}[\frac{1}{2}]$ one has $d^{\pm}t(x+\sum\limits_{j\in S}\frac{1}{2^{j}},0)=d^{\pm}(\sum\limits_{j\in S}\frac{1}{2^{j}},0)=1$, as addition of $x$ changes only finitely many digits. Now assume $x\notin\mathbb{Z}[\frac{1}{2}]$.
[…]
$d^{\pm}(x+\sum\limits_{j\in S_{2}}\frac{1}{2^{j}},0)={\lim\limits_{k\rightarrow\infty}}^{\pm}\frac{\#\{1\leq i\leq k:\,i\notin Q\sqcup S_{2}\}}{k}=$
${\lim\limits_{k\rightarrow\infty}}^{\pm}\Big{(}\frac{\#\{1\leq i\leq k:\,i\notin Q\}}{k}-\frac{\#\{1\leq i\leq k:\,i\in S_{2}\}}{k}\Big{)}=$
${\lim\limits_{k\rightarrow\infty}}^{\pm}\frac{\#\{1\leq i\leq k:\,i\notin Q\}}{k}-\lim\limits_{k\rightarrow\infty}\frac{\#\{1\leq i\leq k:\,i\in S_{2}\}}{k}=$
$={\lim\limits_{k\rightarrow\infty}}^{\pm}\frac{\#\{1\leq i\leq k:\,i\notin Q\}}{k}-0=d^{\pm}(x,0)=c.$
Now it remains to prove the statement only for $S$ such that $S\subseteq Q$. Let us define $R\subseteq\mathbb{N}$ by $\sum\limits_{j\in Q}\frac{1}{2^{j}}+\sum\limits_{j\in S}\frac{1}{2^{j}}=\sum\limits_{j\in R}\frac{1}{2^{j}}$. It is easy to check that:
  • Suppose $i\notin Q$ and $i+1\notin Q$. Then $i\notin R$;
  • Suppose $a\notin Q,\,a+b\notin Q$ and $i\in Q$ for all $a<i<a+b$. Then #$\#\{a\leq i<a+b:\,i\in R\}\leq b-1=\#\{a\leq i<a+b:\,i\in Q\}$.
This immediately implies that $d^{\pm}(x+\sum\limits_{j\in S}\frac{1}{2^{j}},0)\geq d^{\pm}(x,0)=c$ and completes the proof. ∎
Let us also recall simple properties of digit frequencies:
  • $d^{\pm}(2^{k}x,i)=d^{\pm}(2^{l}+x,i)$ for any integers $k,l$;
  • $d^{\pm}(x,i)+d^{\mp}(x,1-i)=1$;
  • If $x\notin\mathbb{Z}[\frac{1}{2}]$, then $d^{\pm}(x,i)=d^{\pm}(-x,1-i)$.
We now focus on the special case $c=\frac{1}{2}$. We will use the construction similar to the set $\mathscr{D}$ but connected to digit frequencies.
Let $\mathscr{F}=D^{-}_{\frac{1}{2}}\cup(\mathscr{D}(\xi)\cap\{x:\,d^{+}(x,0)\geq\frac{1}{2}\})$ where $\xi=\sum\limits_{i=1}^{\infty}\frac{1}{2^{2^{i}}}$. We want to show that the set $\mathscr{F}$ satisfies the conditions of Theorem 3.10 for $m=2$.
Proposition 5.5.
  1. (1)
    $2^{k}\mathscr{F}=2^{k}+\mathscr{F}=\mathscr{F}$ for any integer $k$;
  2. (2)
    $\mathscr{F}\sqcup(-\mathscr{F})\sqcup Q=\mathbb{R}$ where $Q\subseteq\mathbb{Q}$;
  3. (3)
    $\mathscr{F}+\xi\subseteq\mathscr{F}$.
Proof.
  1. (1)
    Follows from the similar properties of all the sets in construction of $\mathscr{F}$;
  2. (2)
    It’s easy to notice that
    =$-\mathscr{F}=\left\{x\notin\mathbb{Z}\left[\frac{1}{2}\right]:\,d^{+}(x,0)<\frac{1}{2}\right\}\cup\left((-\mathscr{D}(\xi))\cap\left\{x:\,d^{-}(x,0)\leq\frac{1}{2}\right\}\right),$
    which shows that $\mathscr{F}$ and $-\mathscr{F}$ are disjoint. If $x\notin\mathscr{F}\sqcup(-\mathscr{F})$, then either $x\in\mathbb{Z}[\frac{1}{2}]$ or $d^{-}(x,0)\leq\frac{1}{2}\leq d^{+}(x,0)$ and $x\notin\mathscr{D}\sqcup(-\mathscr{D})$, which implies $x\in\mathbb{Q}$.
  3. (3)
    By Lemma 5.4 all the sets in the construction of $\mathscr{F}$ satisfy this condition.
This immediately proves that
  • If $\alpha\geq\beta$, then the set $\mathscr{F}$ is $(\alpha,\beta)$-losing;
  • If $\beta>\alpha$ and $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, then the set $\mathscr{F}$ is $(\alpha,\beta)$-winning.
Let’s also note that $\{x:\,d^{-}(x,0)>\frac{1}{2}\}\subseteq\mathscr{F}\subseteq\{x:\,d^{+}(x,0)\geq\frac{1}{2}\}$, thus
Corollary 5.6.
  1. (1)
    If $\alpha\geq\beta$, then the set $D^{-}_{\frac{1}{2}}=\{x:\,d^{-}(x,0)>\frac{1}{2}\}$ is $(\alpha,\beta)$-losing;
  2. (2)
    If $\beta>\alpha$ and $\log_{2}(\alpha\beta)\notin\mathbb{Q}$, then the set $\{x:\,d^{+}(x,0)\geq\frac{1}{2}\}$ is $(\alpha,\beta)$-winning.
Conjecture 5.7.
We believe that $D(D^{-}_{c})=\{(\alpha,\beta):\,\log(\alpha)/\log(\alpha\beta)>c\}$.
6. B-Badly Approximable Numbers
The set of badly approximable numbers is an important object of study in Diophantine approximations. It is defined as follows:
$\operatorname{BA}:=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{n}\right|>\frac{c}{n^{2}}\text{ for some $c>0$ and all }m\in{\mathbb{Z}},\ n\in{\mathbb{N}}\right\}.$
It is known that the badly approximable numbers are dense and have measure zero. Interestingly enough, despite having measure zero, Schmidt showed that they are winning for the largest possible nontrivial bounds:
Theorem 6.1.
[S66] The set $\operatorname{BA}$ is winning for all $(\alpha,\beta)\in\hat{D}$; that is, $D(\operatorname{BA})=\hat{D}$.
In a similar vein to the badly approximable numbers investigated by Schmidt, we define the set of B-Badly Approximable Numbers as
$\operatorname{B-BA}:=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{B^{k}}\right|>\frac{c}{B^{k}}\text{ for some $c>0$ and all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\right\}$
and a closely related set
$\operatorname{B-BA}(c):=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{B^{k}}\right|>\frac{c}{B^{k}}\text{ for all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\right\}.$
These definitions imply $\operatorname{B-BA}=\bigcup_{c>0}\operatorname{B-BA}(c)$. We note that the $\operatorname{B-BA}$ set has been previously studied in [N10] and [PS08].
The following statement follows from work of Schmidt:
Theorem 6.2.
[S66] The set $\operatorname{2-BA}$ is winning for all $(\alpha,\beta)\in\hat{D}$; that is, $D(\operatorname{2-BA})=\hat{D}$.
While both $\operatorname{2-BA}$ and $\operatorname{2-BA}(c)$ did not produce any new Schmidt diagrams, there is a simple change we can make that will produce both nontrivial winning and nontrivial losing zones. For any integer $B\geq 2$, define
$\operatorname{\widetilde{B-BA}}(c,N):=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{B^{k}}\right|>\frac{c}{B^{k}}\text{ for all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\ \text{s.t.}\ k>N\right\}$
and
$\operatorname{\widetilde{B-BA}}(c):=\bigcup_{N\in{\mathbb{N}}}\operatorname{\widetilde{B-BA}}(c,N).$
Throughout this chapter, we define $\gamma_{A}=1-2\alpha+\alpha\beta$ and $\gamma_{B}=1-2\beta+\alpha\beta$. We begin results in this section with a useful general lemma:
Lemma 6.3.
Suppose Bob plays any interval $B_{t}$ and $\gamma_{A}>0$. Denote the center of $B_{t}$ by $c_{t}$. Then, Bob can always choose an interval $B_{t+1}$ such that $|c_{t+1}-c_{t}|\leq\frac{|B_{t}|}{2}\gamma_{A}$. As a corollary, Bob can guarantee that for any $m\in\mathbb{N}$, $|c_{t+m}-c_{t}|\leq\frac{|B_{t}|}{2(1-\alpha\beta)}\gamma_{A}$. Analogous results hold for Alice, with $\gamma_{A}$ replaced by $\gamma_{B}$.
Proof.
We prove Bob’s case; Alice’s case is analogous.
Let’s denote center of $B_{t}$ by $c_{t}$. It is clear that Bob can always choose such a $B_{t+1}$ that $|c_{r+1}-c_{r}|\leq\frac{|B_{t}|}{2}\gamma_{A}$, since no matter which side Alice picks her interval on, Bob will go the opposite way. Extending this logic, it follows that $|c_{m+r}-(a_{0}+k\delta)|\leq\frac{\rho(\alpha\beta)^{m}}{2}\gamma_{A}(1+(\alpha\beta)+(\alpha\beta)^{2}+\ldots)=\frac{\rho(\alpha\beta)^{m}}{2(1-\alpha\beta)}\gamma_{A}$, which completes the proof.
The following theorems provide the winning and losing conditions for $\operatorname{\widetilde{B-BA}}$ sets.
Theorem 6.4.
If
$\frac{2c(1-\beta)(1-\alpha\beta)}{\beta\gamma_{A}}>1,$
then for any $B$, the set $\operatorname{\widetilde{B-BA}}(c)$ is $(\alpha,\beta)$-losing.
Proof.
One has
$\operatorname{\widetilde{B-BA}}(c)^{c}=\left\{x:\,\,\forall N\,\,\exists\,n>N\,\,\text{and}\,\,k\,\in\,\mathbb{Z}:\,\,\left|x-\frac{k}{B^{n}}\right|<\frac{c}{B^{n}}\right\}=$
$=\bigcap\limits_{N=1}^{\infty}\bigcup\limits_{n=N}^{\infty}\left(\bigcup\limits_{k\in\mathbb{Z}}\left[\frac{k}{B^{n}}-\frac{c}{B^{n}},\frac{k}{B^{n}}+\frac{c}{B^{n}}\right]\right)\subseteq\bigcap\limits_{N=1}^{\infty}\bigcup\limits_{n=N}^{\infty}\left(\bigcup\limits_{k\in\mathbb{Z}}\left(\frac{k}{B^{n}}-\frac{c}{B^{n}},\frac{k}{B^{n}}+\frac{c}{B^{n}}\right)\right),$
and let $J_{n}=\bigcup\limits_{k\in\mathbb{Z}}\left(\frac{k}{B^{n}}-\frac{c}{B^{n}},\frac{k}{B^{n}}+\frac{c}{B^{n}}\right)$. Fix $\varepsilon>0$ such that
$\varepsilon=\frac{2c(1-\beta)(1-\alpha\beta)}{\beta\gamma_{A}}-1.$
Bob will play in the following way:
  • On his first turn, Bob will play an interval of length
    $\frac{1}{B}\left(\frac{2c(1-\alpha\beta)}{\gamma_{A}}\right)$
    centered at $1/B$. By Lemma 1.2, Bob can play in such a way that in finitely many moves, $B_{s}\subset J_{1}$.
  • On all subsequent turns, Bob plays arbitrarily until some turn $t$ such that
    $\log_{B}\left(\frac{1}{1-\beta}\right)$$\leq\left\{\left(t\cdot\log_{B}\alpha\beta\right)+\log_{B}|A_{0}|\right\}$ (6.1)
    $\leq\log_{B}\left(\frac{1}{1-\beta}\right)+\log_{B}(1+\varepsilon),$
    there exists $N\in\mathbb{N}$ such that
    $\left(\frac{1}{1-\beta}\right)\leq B^{N}|A_{t}|\leq\left(\frac{1}{1-\beta}\right)(1+\varepsilon)=\left(\frac{2c(1-\alpha\beta)}{\beta\gamma_{A}}\right),$ (6.2)
    and since $|A_{t}|(1-\beta)>1/B^{N}$, Bob can play his next interval centered at $m/B^{N}$ for some $m\in\mathbb{Z}$. Furthermore, by Lemma 1.2, Bob can play in such a way that in finitely many moves, $B_{t+s}\subset J_{n}$. By the Equidistribution Theorem, there exists an infinite number of possible $t$ that satisfy inequality 6.2. Therefore, Bob can do this an infinite number of times, and Alice loses.
In the special case of the $\operatorname{\widetilde{2-BA}}$ set, we have the following theorem.
Theorem 6.5.
The set $\operatorname{\widetilde{2-BA}}(c)$ is winning for $\alpha<\frac{1}{3}$, $\beta>\frac{9c}{1-3\alpha}$, and $c<1/6$.
Proof.
The winning strategy for Alice is as follows: suppose Alice is given an interval $B_{i}$ of length $l_{i}$. Pick $k_{i}$ such that
$\frac{1}{2^{k_{i}}}<l_{i}\leq\frac{1}{2^{k_{i}-1}}$
or equivalently,
$\frac{l_{i}}{2}\leq\frac{1}{2^{k_{i}}}<l_{i}.$
Let $W_{i}$ denote the set of all intervals of length $\alpha l_{i}$ and center $c_{i}$ of the form $(m\pm\frac{1}{3})2^{-k_{i}}$, where $m\in{\mathbb{Z}}$. Then, Alice will select an interval from $W_{i}$. This is always possible because the maximum distance between the two endpoints of sets is
$\left|c_{i}-c^{\prime}_{i}\right|+\alpha l_{i}=\frac{1}{3\cdot 2^{k_{i}-1}}+\alpha l_{i}<\frac{2l_{i}}{3}+\frac{l_{i}}{3}=l_{i}$
and thus any interval of length $l_{i}$ must contain at least one of the intervals in $W_{i}$.
Next, define $q_{i}$ to be the smallest integer such that $m/2^{q_{i}}\in A_{i}$ for some $m\in{\mathbb{Z}}$. Note that
$\left|\frac{m}{2^{q_{i}}}-c_{i}\right|=\frac{1}{3\cdot 2^{q_{i}}}<\frac{\alpha l_{i}}{2}<\frac{1}{3\cdot 2^{q_{i}-1}}.$
We wish to show that for all $i\geq 1$ and $x\in A_{i}$,
$\left|x-\frac{m}{2^{n}}\right|>\frac{c}{2^{n}}\ \text{for all}\ n\leq q_{i}-2.$
by induction on $i$. Suppose that the statement holds for $i$. Then, Bob picks an interval $B_{i+1}$ with length $l_{i+1}=\alpha\beta l_{i}$. Now, note that for all $x\in A_{i+1}$ and $q_{i}\leq n\leq k_{i+1}$,
$\left|x-\frac{m}{2^{n}}\right|\geq\left|x-\frac{m^{\prime}}{2^{k_{i+1}}}\right|$ $>\frac{1}{3\cdot 2^{k_{i+1}}}-\frac{\alpha l_{i+1}}{2}$
$\geq\frac{l_{i+1}}{6}-\frac{\alpha l_{i+1}}{2}$
$=\alpha\beta l_{i}\left(\frac{1-3\alpha}{6}\right)$
$>\alpha l_{i}\left(\frac{9c}{1-3\alpha}\right)\left(\frac{1-3\alpha}{6}\right)$
$=\frac{3}{2}\alpha l_{i}c$
$>\frac{c}{2^{q_{i}}}\geq\frac{c}{2^{n}}.$
To take care of $k_{i+1}+1\leq n\leq q_{i+1}-2$, simply note that for all $x\in A_{i+1}$ and $c<1/6$,
$\left|x-\frac{m}{2^{n}}\right|$ $>\frac{1}{3\cdot 2^{n}}-\frac{\alpha l_{i+1}}{2}>\frac{1}{3\cdot 2^{n}}-\frac{1}{3\cdot 2^{q_{i+1}-1}}>\frac{c}{2^{n}}.$
Finally, for $n=q_{i}-1$, we split into cases based on $A_{i}$, the prior move.
  • Case I.
    If
    $\frac{\alpha l_{i}}{2}<\frac{1}{3\cdot 2^{q_{i}-1}}-\frac{c}{2^{q_{i}-1}},$
    then for all $x\in A_{i}$,
    $\left|x-\frac{m}{2^{q_{i}-1}}\right|$ $\geq\frac{1}{3\cdot 2^{q_{i}-1}}-\frac{\alpha l_{i}}{2}$
    $>\frac{1}{3\cdot 2^{q_{i}-1}}-\frac{1}{3\cdot 2^{q_{i}-1}}+\frac{c}{2^{q_{i}-1}}$
    $=\frac{c}{2^{q_{i}-1}}.$
  • Case II.
    Assume
    $\frac{\alpha l_{i}}{2}\geq\frac{1}{3\cdot 2^{q_{i}-1}}-\frac{c}{2^{q_{i}-1}}.$
    Then, if
    $B_{i+1}\cap\left[\frac{m-c}{2^{q_{i}-1}},\frac{m+c}{2^{q_{i}-1}}\right]=\emptyset,$
    it is clear that Alice does not have to worry about $n=q_{i}-1$, since for all $x\in A_{i+1}\subset B_{i+1}$, $|x-m/2^{q_{i}-1}|>c/2^{q_{i}+1}$. Thus, for the remainder of this case, assume $B_{i+1}$ intersects the interval. Without loss of generality, assume that $m/2^{q_{i}-1}>x$ for all $x\in B_{i+1}$ (that is, the fraction lies to the right of the interval $B_{i+1}$). This is possible because the mapping $x\mapsto-x$ gives us the other case. Alice will then choose an interval $A_{i+1}\in W_{i+1}$ centered at $c_{i+1}$ such that
    $c_{i+1}\leq\frac{m}{2^{q_{i}-1}}-\frac{2}{3}\left(\frac{1}{2^{k_{i+1}}}\right),$
    that is, Alice cannot choose the closest interval to $\frac{m}{2^{q_{i}-1}}$. This is always possible because the second closest interval is within $l_{i+1}$ of $\frac{m}{2^{q_{i}-1}}$; that is, for $x\in A_{i+1}$,
    $\left|x-\frac{m}{2^{q_{i}-1}}\right|<\frac{2}{3}\left(\frac{1}{2^{k_{i+1}}}\right)+\frac{\alpha l_{i+1}}{2}<\frac{2}{3}l_{i+1}+\frac{1}{6}l_{i+1}<l_{i+1}$
    and thus whenever Alice is able to pick the closest interval, Alice is also able to pick the second closest interval. Furthermore, for $x\in A_{i+1}$,
    $\left|x-\frac{m}{2^{q_{i}-1}}\right|$ $>\frac{2}{3}\left(\frac{1}{2^{k_{i+1}}}\right)-\frac{\alpha l_{i+1}}{2}$
    $>\frac{l_{i+1}}{3}-\alpha l_{i+1}$
    $=\alpha\beta l_{i}\left(\frac{1-3\alpha}{3}\right)$
    $>3\alpha l_{i}c$
    $>\frac{c}{2^{q_{i}-1}}$
    proving this case and completing the induction.
Thus, $x_{\infty}\in\operatorname{\widetilde{2-BA}}(c)$ and Alice wins.
A special case of winning and losing zones given by Theorem 6.4 and Theorem 6.5 can be seen on Figure 5.
We also consider the special case of $\widetilde{\text{3-BA}}(1/6)$, defined as
$\widetilde{\text{3-BA}}(1/6):=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{3^{k}}\right|\geq\frac{1}{6\cdot 3^{k}}\text{ for all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\right\}.$
There, we have some additional bounds:
Theorem 6.6.
If
$\beta\gamma_{A}<\gamma_{B}$
and
$\frac{1-\alpha\beta}{\alpha\gamma_{B}}\geq\frac{9\beta}{1-\beta}$
and
$\alpha\beta>\frac{1}{7},$
then the set $\widetilde{\text{3-BA}}(1/6)$ is $(\alpha,\beta)$-losing.
Proof.
Let
$2\varepsilon={\gamma_{B}}-\beta\gamma_{A}.$
Then, Bob will play in the following way: on his first turn, he plays an interval of length
$|B_{0}|=\frac{1-\alpha\beta}{\alpha\gamma_{B}}.$
Then, he will play arbitrarily until some turn $t$, where
$\frac{1-\alpha\beta}{\alpha\gamma_{B}}\cdot\frac{1}{3^{n+2}}\leq|B_{t}|<(1+\varepsilon)\cdot\frac{1-\alpha\beta}{\alpha\gamma_{B}}\cdot\frac{1}{3^{n+2}}<\frac{1-\alpha\beta}{\alpha\beta\gamma_{A}}\cdot\frac{1}{3^{n+2}}$
for some $n\in\mathbb{N}$. Then, on turn $t$, he will center his ball at a $c_{t}$ such that
$c_{t}=\frac{m}{3^{n}}$
for some $m\in\mathbb{Z}$. On his next move, Bob is always able to play in one of the following two ways:
  • Bob will center his next interval such that
    $\left|c_{t+1}-\frac{m}{3^{n}}\right|<\frac{1}{6\cdot 3^{n}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$ (6.3)
    and Bob will try to minimize $|c_{t+1}-\frac{m}{3^{n}}|$, or
  • Bob will center his next interval such that
    $\left|c_{t+1}-\left(\frac{m}{3^{n}}\pm\frac{1}{3^{n+1}}\right)\right|<\frac{1}{6\cdot 3^{n+1}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$ (6.4)
    and Bob will try to minimize $|c_{t+1}-\frac{m}{3^{n}}\mp\frac{1}{3^{n}+1}|$.
We show that this is always possible. First, we show that Bob can play his strategy on turn $t$. The length condition is equivalent to
$0\leq\left\{\log_{3}\frac{|B_{0}|\alpha\gamma_{B}}{1-\alpha\beta}+t\log_{3}\alpha\beta\right\}<\log_{3}(1+\varepsilon)$
and by the Equidistribution Theorem, there are infinitely many values of $t$ that satisfy the inequality. Then, we have
$|B_{t}|\geq\frac{1-\alpha\beta}{\alpha\beta\gamma_{A}}\cdot\frac{1}{3^{n+2}}\geq\frac{\beta}{1-\beta}\cdot\frac{1}{3^{n}}$
so $(1-\beta)|A_{t-1}|\geq 1/3^{n}$ and therefore Bob can always play an interval of the form $m/3^{n}$ on turn $t$.
Now, we show Bob can play his strategy on turn $t+1$. First, we note that since $\alpha\beta>1/7$, we have
$\frac{1}{6\alpha\beta}$ $<\frac{7}{6}$
$\frac{|B_{t}|\gamma_{A}}{2(1-\alpha\beta)}<\frac{1}{2\alpha\beta}\cdot\frac{1}{3^{n+2}}$ $<\frac{7}{6\cdot 3^{n+1}}$
$\frac{|B_{t}|\gamma_{A}}{2}+\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$ $<\frac{1}{3^{n+1}}+\frac{1}{6\cdot 3^{n+1}}$
$\frac{m}{3^{n}}+\frac{|B_{t}|\gamma_{A}}{2}$ $<\frac{m}{3^{n}}+\frac{1}{3^{n+1}}+\frac{1}{6\cdot 3^{n+1}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)},$
which implies that Bob is able to ensure that $c_{t+1}<\frac{m}{3^{n}}+\frac{1}{3^{n+1}}+\frac{1}{6\cdot 3^{n+1}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$. By similar logic, Bob can ensure that $c_{t+1}>\frac{m}{3^{n}}-\frac{1}{3^{n+1}}-\frac{1}{6\cdot 3^{n+1}}+\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$. Furthermore, we note that
$\frac{|B_{t}|\alpha\gamma_{B}}{1-\alpha\beta}$ $>\frac{1}{3^{n+2}}$
$|B_{t}|\alpha\left(1-\beta-\frac{\beta\gamma_{A}}{1-\alpha\beta}\right)$ $>\frac{1}{3^{n+2}}$
$\frac{1}{3^{n+2}}+\frac{|B_{t}|\alpha\beta\gamma_{A}}{1-\alpha\beta}$ $<(1-\beta)\alpha|B_{t}|$
$\left|\left(\frac{m}{3^{n}}\pm\left(\frac{1-\frac{1}{6}}{3^{n+1}}+\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}\right)\right)-\left(\frac{m}{3^{n}}\pm\left(\frac{1}{6\cdot 3^{n}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}\right)\right)\right|$ $<(1-\beta)\alpha|B_{t}|,$
which implies that Bob is always able to satisfy either 6.3 or 6.4.
Finally, we show this strategy is winning. This follows from Lemma 6.3: if 6.3 is true, then in finitely many $k$ turns, $|B_{t+k}|\subset(\frac{m-1/6}{3^{n}},\frac{m+1/6}{3^{n}})$. Since Bob can play this way for infinitely many sets of two consecutive turns, Bob wins. ∎
Conjecture 6.7.
If
$\beta\gamma_{A}\geq\gamma_{B},$
then the set $\widetilde{\text{3-BA}}(1/6)$ is $(\alpha,\beta)$-winning. If not, then it is $(\alpha,\beta)$-losing.
Acknowledgements
We are very grateful to Prof. Dmitry Kleinbock for the project proposal and his valuable suggestions in research paths. We would also like to thank Prof. Pavel Etingof, Dr. Slava Gerovitch, Dr. Felix Gotti, Dr. Tanya Khovanova, and all the PRIMES organizers for providing the opportunity to conduct math research through this program.

### v3

@@
 Abstract.
-In this paper we completely describe the winning and losing conditions different from the only “trivial” conditions known before.
-In other words, we solve the open question of finding a complete nontrivial Schmidt diagram.
-In addition, we give the new bounds for two family of sets: one related to frequencies of digits in base-$2$ expansions, and one connected to the set of the badly approximable numbers.
+In this paper we study the classical Schmidt game on two families of sets: one related to frequencies of digits in base-$2$ expansions, and one connected to the set of the badly approximable numbers.
+Namely, we describe some nontrivial winning and losing parameters $(\alpha,\beta)$ for these sets.
 1. Set-up and Introduction: Schmidt Games
 Wolfgang Schmidt first introduced Schmidt games in [S66] in order to study the set of badly approximable numbers.
-Winning sets in these games have special properties, such as a full Hausdorff dimension; as such, Schmidt games have been used to prove results in number theory and dynamical systems.
-Here, we briefly describe the setup of a Schmidt $(\alpha,\beta)$-game.
+These games involve two players, Alice and Bob and two parameters $\alpha,\beta\in(0,1)$.
+A subset $S$ of a metric space is called $(\alpha,\beta)$-winning if Alice has a strategy guaranteeing that a certain point, constructed as a result of the infinite series of moves, lies in $S$.
+Of particular interest are sets that are $(\alpha,\beta)$-winning for all $\beta$ and a fixed $\alpha$; those are called $\alpha$-winning sets.
+Schmidt proved that $\alpha$-winning sets have full Hausdorff dimension, and a countable intersection of $\alpha$-winning sets is $\alpha$-winning.
+The technique of Schmidt games has been frequently used to prove that certain sets arising in number theory and dynamical systems have full Hausdorff dimension.
+Many of the applications of Schmidt games deal with various stronger modifications of Schmidt’s original construction.
+These modifications include the McMullen’s strong and absolute winning games [M10], with an additional property that the winning property is preserved by quasisymmetric homeomorphisms; hyperplane absolute winning game [BFKRW12], where the winning property holds under $C^{1}$ diffeomorphisms; potential game [FSU18], Cantor game [BH17], and Hausdorff/packing games [DFSU24] which give a tool to precisely calculate Hausdorff/packing dimensions.
+For a detailed survey describing the setups of some of the aforementioned games and connections between them, see [BHNS25].
+However, in this paper, we restrict our attention to the $(\alpha,\beta)$-game introduced in [S66].
+Here, we briefly describe the setup of the classical Schmidt $(\alpha,\beta)$-game.
 Throughout this paper we work with the Schmidt game on $\mathbb{R}$, so for simplicity we give the definitions only in this special case.
@@
 Then, fix a set $S\subset{\mathbb{R}}$ and $(\alpha,\beta)\in I$.
-Two players, Alice and Bob, can play a Schmidt game on $S$, taking alternating turns, as follows: Bob begins by choosing an interval $B_{0}$.
-Given an interval $B_{i}$, Alice chooses an interval $A_{i}\subset B_{i}$ such that $|A_{i}|=\alpha|B_{i}|$.
+Two players, Alice and Bob, can play a Schmidt game on $S$, taking alternating turns, as follows: Bob begins by choosing a closed interval $B_{0}$.
+Given an interval $B_{i}$, Alice chooses a closed interval $A_{i}\subset B_{i}$ such that $|A_{i}|=\alpha|B_{i}|$.
 Similarly, given an interval $A_{i}$, Bob chooses an interval $B_{i+1}\subset A_{i}$ such that $|B_{i+1}|=\beta|A_{i}|$.
@@
 We call $S$ $(\alpha,\beta)$-winning if, regardless of how Bob chooses his intervals, Alice can play in such a way that $x_{\infty}\in S$.
-For every $S$, define the Schmidt diagram $D(S)$ of $S$ as the set of all pairs $(\alpha,\beta)\in I$ such that $S$ is $(\alpha,\beta)$-winning.
-In his original paper, Schmidt proved that there are two “trivial zones” of Schmidt diagrams:
+In this paper, we will often fix the set $S$ and explore the games for different pairs $(\alpha,\beta)$. To do this, we will utilize a little bit different language.
+Let’s fix the set $S$.
+A point $(\alpha,\beta)\in I$ is called winning for $S$ if $S$ is $(\alpha,\beta)$-winning and non-winning for $S$ otherwise.
+A point $(\alpha,\beta)\in I$ is called losing for $S$ if, regardless of how Alice chooses her intervals, Bob can play in such a way that $x_{\infty}\notin S$, and non-losing otherwise.
+It will be crucial for some of our proofs to know that for a given set $S$ and any parameters $(\alpha,\beta)$ one of the players always has a strategy to win; that is, we will require that all the non-losing parameters are winning, and all the non-winning parameters are losing.
+In this case, we will say that the Schmidt game is determined on the set $S$.
+In general, Schmidt game can be not determined: see [AFG23] for some constructions using the Axiom of choice.
+However, it is known ([CFJ23] ) that the Schmidt game in $\mathbb{R}$ is determined if we assume the Axiom of determinacy AD (which contradicts the Axiom of choice).
+More specifically, it is known that the Schmidt game is determined on Borel sets.
+For every $S$, define the Schmidt diagram $D(S)$ of $S$ as the set of all winning $(\alpha,\beta)\in I$. In his original paper, Schmidt proved that there are two “trivial zones” of Schmidt diagrams:
 Lemma 1.1.
@@
 Figure 1. $\check{D}$ (blue)
-Figure 1. $\check{D}$ (blue)
 Figure 2. $\hat{D}$ (blue)
-There are only four sets which were known to coincide with $D(S)$ for some $S$: $\emptyset$, $\check{D}$, $\hat{D}$, and $I$.
-Our goal with this project was to find some set $S$ such that $\check{D}\subsetneq D(S)\subsetneq\hat{D}$ and to completely describe $D(S)$.
-It is worth noting that this problem was investigated in an earlier PRIMES paper, which investigated sets relating to particular decimal expansions [KZT12].
-In the aforementioned paper, some nontrivial statements about Schmidt diagrams were obtained, but no $D(S)$ different from the four diagrams discussed above was completely described.
+There are only four sets that are known to coincide with $D(S)$ for some $S$: $\emptyset$, $\check{D}$, $\hat{D}$, and $I$.
+It is clear that this is far not the complete list of the possible Schmidt diagrams, see [F82] for an example of the set which is proven to have a nontrivial diagram, however, no diagram $D(S)$ different from the four diagrams discussed above was completely described.
+It is worth noting that the problem of completely describing a nontrivial Schmidt diagram was investigated in an earlier PRIMES paper [KZT12], which investigated sets relating to particular decimal expansions.
+In this paper, we partially describe the nontrivial Schmidt diagrams of two families of sets, and make some conjectures about their precise form.
+The structure of the paper is as follows.
+In Section 2, we formulate our main results.
+We describe some winning and losing conditions for families of sets connected with digit frequencies in the dyadic expansion of real numbers in Subsection 2.1, and for families of sets connected with badly approximable numbers in base $B$ expansion in Subsection 2.2.
+In Section 3, we discuss some general properties of winning sets and general symmetric strategies on special families of sets, which we need to prove Theorem 2.2.
+We prove Theorem 2.2 in Section 4 and the theorems from Subsection 2.2 in Section 5.
 2. Main results
-One of the main results of this paper is a complete description of a Schmidt Diagram different from $\emptyset$, $\check{D}$, $\hat{D}$ and $I$. More specifically, we prove the following theorem.
+2.1. Digit frequencies and related sets
+We will investigate a construction regarding frequencies of digits in base-$2$ expansions of real numbers. These sets have been studied in the past (see [V59, FFNS21]).
+First, some setup.
+Consider the base-$2$ expansions of the form $x=x_{0},x_{1}x_{2}\cdots$ where $x_{0}$ is an integer and $x_{i}\in\{0,1\}$ are the digits in the base-$2$ expansion of $x$.
+In this paper, if $x$ can be represented by two separate expansions, we take the lexicographically minimal expansion (for example, $0.100\ldots$ would instead be written as $0.011\ldots$).
+We define
+$d^{+}(x,j)=\limsup_{k\rightarrow\infty}\frac{\#\{1\leq i\leq k:x_{i}=j\}}{k}$
+and similarly,
+$d^{-}(x,j)=\liminf_{k\rightarrow\infty}\frac{\#\{1\leq i\leq k:x_{i}=j\}}{k}.$
+Define the sets
+$F^{+}_{c}$ :=$:=\{x\in{\mathbb{R}}:d^{+}(x,0)>c\},$
+$F^{-}_{c}$ :=$:=\{x\in{\mathbb{R}}:d^{-}(x,0)>c\}.$
+It is known ([Bo09]) that the set $\{x\in\mathbb{R}:\,\,d^{-}(x,j)=d^{+}(x,j)=\frac{1}{2}\}$ has full Lebesgue measure for any $j\in\{0,1\}$.
+In particular, the sets $F^{\pm}_{c}$ have full Lebesgue measure if $c<\frac{1}{2}$ and measure zero otherwise.
+Moreover, Besicovitch ([B35]; see also [O04] for generalizations) showed that the Hausdorff dimension of $F^{-}_{c}$ for $c>1$ is equal to $-c\log_{2}c-(1-c)\log_{2}(1-c)$.
+Winning and losing parameters of these sets were studied before. The following fact is not stated, but follows from the proof of Theorem 2.15 in [FFNS21].
+Proposition 2.1.
+[FFNS21] Let $c\in(0,1)$. The set $F^{+}_{c}$ is winning for $\log(8\alpha)/\log(\alpha\beta)>c$ and losing for $\log(\alpha/8)/\log(\alpha\beta)\leq c$.
+The results of our research expand on that bound:
 Theorem 2.2.
-There exists a set $\mathscr{S}$ such that $D(\mathscr{S})=\{(\alpha,\beta):0<\alpha<\beta<1\}$ (see Figure 3).
-Figure 3. $D(\mathscr{S})$ (blue)
-We develop the techniques for this in Section 3, and the more general statements about sets with such a diagram are given in Theorem 3.10 and Corollary 3.11.
-We construct a special family of examples in Section 4.
-In Section 5, we describe some winning and losing conditions for families of sets connected with digit frequencies in the dyadic expansion of real numbers.
-We prove a general case in Theorem 5.3 and a special symmetric case in Corollary 5.6.
-In Section 6 we give some new results about winning and losing conditions for families of sets connected with badly approximable numbers in base $B$ expansion: specifically in Theorems 6.4, 6.5, and 6.6.
-The bounds for the winning and losing zones we describe in Sections 5 and 6 are not optimal; however, they give some estimates and shed some light on how their respective Schmidt diagrams can look like.
-Figures 5 and 5 show two examples of the possible bounds (where the blue zones are subsets of the corresponding Schmidt diagram and the red zones are subsets of its complement).
-The sets $D_{c}^{-}$ used in the first example will be defined later in Section 5 and the sets $\operatorname{\widetilde{2-BA}}(c)$ will be defined in Section 6.
-Figure 4. Bounds for $D(D_{c}^{-})$ with $c=0.7$
-Figure 4. Bounds for $D(D_{c}^{-})$ with $c=0.7$
+Let $c\in(0,1)$, and suppose $\log_{2}{\alpha\beta}\notin\mathbb{Q}$.
+Then, the sets $F^{\pm}_{c}$ are $(\alpha,\beta)$-winning if $\log(4\alpha)/\log(\alpha\beta)>c$, and $(\alpha,\beta)$-losing if $\log(\alpha/4)/\log(\alpha\beta)\leq c$.
+In addition:
+  1. (1)
+    If $c<\frac{1}{2}$, then the set $F^{+}_{c}$ is $(\alpha,\beta)$-winning for any $(\alpha,\beta)$ such that $\beta>\alpha$ and $\log_{2}(\alpha\beta)\notin\mathbb{Q}$.
+  2. (2)
+    If $c\geq\frac{1}{2}$, then the set $F^{-}_{c}$ is $(\alpha,\beta)$-losing for any $(\alpha,\beta)$ such that $alpha\geq\beta$.
+Here we do not require the restriction $\log_{2}{\alpha\beta}\notin\mathbb{Q}$.
+Let us also notice that Schmidt’s original approach (see Corollary 1 of Theorem 6 in [S66]) automatically gives us some losing zone bounds as a function of Hausdorff dimension.
+Namely, for any set $S\subseteq\mathbb{R}$ the following statement holds: if $\dim_{H}S=s$, and $\frac{\log\lfloor\beta^{-1}\rfloor}{-\log(\alpha\beta)}>s$, then $S$ is $(\alpha,\beta)$-non-winning.
+Using the fact that $\dim_{H}F_{c}^{-}=-c\log_{2}c-(1-c)\log_{2}(1-c)$ for $c>\frac{1}{2}$, we automatically get the following statement:
+Proposition 2.3.
+Let $c>\frac{1}{2}$, and suppose $\frac{\log\lfloor\beta^{-1}\rfloor}{-\log(\alpha\beta)}>-c\log_{2}c-(1-c)\log_{2}(1-c)$. Then, the set $F_{c}^{-}$ is $(\alpha,\beta)$-losing.
+The winning and losing zones for selected values of $c$ are shown at figures  4,  4,  6 and  6.
+The winning zones from Theorem 2.2 are in blue, the losing zones from Theorem 2.2 are in red and the losing zones from Proposition 2.3 are in green.
+We note that the losing conditions in Theorem 2.2 are always stronger that the ones from Proposition 2.3 for small enough $\alpha$.
+Figure 3. Bounds for $D(F^{+}_{c})$ with $c=0.1$
+Figure 4. Bounds for $D(F^{+}_{c})$ with $c=0.55$
+Figure 5. Bounds for $D(F^{-}_{c})$ with $c=0.9$
+Figure 6. Bounds for $D(F^{-}_{c})$ with $c=0.45$
+We will develop the necessary techniques and prove Theorem 2.2 in Section 3. Here, we would like to state a conjecture which our methods and results lead to.
+Conjecture 2.4.
+We believe that $D(F^{\pm}_{c})=\{(\alpha,\beta):\,\log(\alpha)/\log(\alpha\beta)>c\}$. In particular, we expect that $D(F^{\pm}_{\frac{1}{2}})=\{(\alpha,\beta):\,\alpha<\beta\}$.
+2.2. d-Badly approximable numbers
+The set of badly approximable numbers is an important object of study in Diophantine approximations. It is defined as follows:
+$\operatorname{BA}:=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{n}\right|>\frac{c}{n^{2}}\text{ for some $c>0$ and all }m\in{\mathbb{Z}},\ n\in{\mathbb{N}}\right\}.$
+It is known that the badly approximable numbers are dense and have Lebesgue measure zero.
+Schmidt showed in [S66] that they are winning for the largest possible bounds: namely, the set $\operatorname{BA}$ is winning for all $(\alpha,\beta)\in\hat{D}$.
+That is, $D(\operatorname{BA})=\hat{D}$; in particular, the set $\operatorname{BA}$ is $\alpha$-winning for $\alpha<\frac{1}{2}$ and $\dim_{H}\operatorname{BA}=1$.
+In a similar vein to the badly approximable numbers investigated by Schmidt, we define the set of d-Badly Approximable Numbers as
+$\operatorname{d-BA}:=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{d^{k}}\right|>\frac{c}{d^{k}}\text{ for some $c>0$ and all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\right\}$
+and a closely related set
+$\operatorname{d-BA}(c):=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{d^{k}}\right|>\frac{c}{d^{k}}\text{ for all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\right\}.$
+These definitions imply $\operatorname{d-BA}=\bigcup_{c>0}\operatorname{d-BA}(c)$.
+It also follows from the work of Schmidt ([S66]) that the set $\operatorname{2-BA}$ is winning for all $(\alpha,\beta)\in\hat{D}$; that is, $D(\operatorname{2-BA})=\hat{D}$.
+While both $\operatorname{2-BA}$ and $\operatorname{2-BA}(c)$ did not produce any new Schmidt diagrams, there is a simple change we can make that will produce both nontrivial winning and nontrivial losing zones.
+For any integer $d\geq 2$, define
+$\operatorname{\widetilde{d-BA}}(c,N):=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{d^{k}}\right|>\frac{c}{d^{k}}\text{ for all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\ \text{s.t.}\ k>N\right\}$
+and
+$\operatorname{\widetilde{d-BA}}(c):=\bigcup_{N\in{\mathbb{N}}}\operatorname{\widetilde{d-BA}}(c,N).$
+We note that these sets has been previously studied in [N10] and [PS08].
+In particular, the Lebesgue measure of all the aforementioned sets is zero, and the precise Hausdorff dimension of $\operatorname{2-BA}$ for any parameter $c$ is calculated in [N10].
+Moreover, there exists such a critical value $c_{0}$ that the set $\operatorname{2-BA}$ is countable if and only if $c>c_{0}$.
+This value is known to be related to Thue-Morse sequences; see Theorem 25 in [N10] for details.
+Let us define $\gamma=1-2\alpha+\alpha\beta$.
+The following theorems provide the winning and losing conditions for $\operatorname{\widetilde{d-BA}}$ sets.
+Theorem 2.5.
+Fix $0<c<\frac{1}{2}$. If
+$\frac{2c(1-\beta)(1-\alpha\beta)}{\beta\gamma}>1,$
+then for any $d$, the set $\operatorname{\widetilde{d-BA}}(c)$ is $(\alpha,\beta)$-losing.
+In the special case of the $\operatorname{\widetilde{2-BA}}$ set, we have the following theorem.
+Theorem 2.6.
+Fix $0<c<\frac{1}{6}$. The set $\operatorname{\widetilde{2-BA}}(c)$ is $(\alpha,\beta)$-winning if $\alpha<\frac{1}{3}$ and $\beta>\frac{9c}{1-3\alpha}$.
+Just as with the sets $F_{c}^{-}$, we note that some losing parameters for the sets $\operatorname{\widetilde{2-BA}}$ can be found using the Hausdorff dimension of $\operatorname{\widetilde{2-BA}}$.
+It is not hard to see that for small enough $c$ these losing zones we can obtain this way are contained in the zones given by Theorem 2.5.
+A special case of winning and losing zones given by Theorem 2.5 and Theorem 2.6 can be seen at Figure 7.
 Figure 7. Bounds for $D(\operatorname{\widetilde{2-BA}}(c))$ with $c=0.035$
-3. General strategies in the Schmidt’s game
+3. General strategies in the Schmidt game
 This section gives some general statements concerning strategies in Schmidt’s game.
+These statements will be used to prove Theorem 2.2; however, we believe they may also prove useful for proving results concerning Schmodt games on a larger variety of sets.
 Let us start with a classical result due to Schmidt:
-Lemma 3.1.
+Proposition 3.1.
 [S66] If a set $S$ is $(\alpha,\beta)$-winning, $\alpha\beta=\alpha^{\prime}\beta^{\prime}$, and $\alpha>\alpha^{\prime}$, then $S$ is also $(\alpha^{\prime},\beta^{\prime})$-winning.
 Sometimes for simplicity we will also refer to existence the of positional strategies, i.e. to the following classical statement:
-Theorem 3.2.
+Proposition 3.2.
 [S66] If a set $S$ is winning, there exists a positional winning strategy; that is, a strategy that only depends on the most recent interval played and not the prior history of moves.
@@
 Fix $m\in\mathbb{N}$.
-Let $D\subseteq\mathbb{R}$ be some set with a property that $m^{k}D=m^{k}+D=D$ for any integer $k$.
-Suppose Bob has a strategy for $(\alpha,\beta)$-game with the set $D$ that allows him to win, and the first ball in this strategy is $B_{0}$.
+Let $S\subseteq\mathbb{R}$ be some set with a property that $m^{k}=m^{k}+S=S$ for any integer $k$.
+Suppose Bob has a strategy for $(\alpha,\beta)$-game with the set $S$ that allows him to win, and the first ball in this strategy is $B_{0}$.
 Then:
   1. (1)
-    For any $k\in\mathbb{Z}$ Bob has a winning strategy for $(\alpha,\beta)$-game with the set $D$ starting with the ball $m^{k}B_{0}$;
+    For any $k\in\mathbb{Z}$ Bob has a winning strategy for $(\alpha,\beta)$-game with the set $S$ starting with the ball $m^{k}B_{0}$;
   2. (2)
-    For any $k\in\mathbb{Z}$ Bob has a winning strategy for $(\alpha,\beta)$-game with the set $D$ starting with the ball $m^{k}+B_{0}$.
+    For any $k\in\mathbb{Z}$ Bob has a winning strategy for $(\alpha,\beta)$-game with the set $S$ starting with the ball $m^{k}+B_{0}$.
 (Here $m^{k}B_{0}=\{m^{k}x:\,\,\,x\in B_{0}\}$ and $m^{k}+B_{0}=\{m^{k}+x:\,\,\,x\in B_{0}\}$).
@@
     It is easy to see that $\bigcap\limits_{i=0}^{\infty}B_{i}^{\prime}=m^{k}\bigcap\limits_{i=0}^{\infty}B_{i}=m^{k}x$, where $x$ is the outcome of some game with the starting ball $B_{0}$, i.e.
-$x\notin D$, therefore $m^{k}x\notin D$ and Bob wins.
+$x\notin S$, therefore $m^{k}x\notin S$ and Bob wins.
   2. (2)
-    The strategy and the argument is analogous; we just take $A_{i}=A_{i}^{\prime}-m^{k}$ and $B_{i}^{\prime}=B_{i}+m^{k}$ and use the fact that $x\in D$ if and only if $m^{k}+x\in D$.
+    The strategy and the argument is analogous; we just take $A_{i}=A_{i}^{\prime}-m^{k}$ and $B_{i}^{\prime}=B_{i}+m^{k}$ and use the fact that $x\in S$ if and only if $m^{k}+x\in S$.
 Lemma 3.4.
@@
 Fix $m\in\mathbb{N}$.
-Let $D\subseteq\mathbb{R}$ be some set with a property that $m^{k}D=m^{k}+D=D$ for any integer $k$.
-Let parameters $(\alpha,\beta)$ be such that $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, and suppose Bob wins $(\alpha,\beta)$-game with $D$.
+Let $S\subseteq\mathbb{R}$ be some set with a property that $m^{k}S=m^{k}+S=S$ for any integer $k$.
+Let parameters $(\alpha,\beta)$ be such that $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, and suppose Bob wins $(\alpha,\beta)$-game with $S$.
 Then for any $L>0$ and any $\varepsilon>0$ there exists a winning strategy in this game for Bob starting with $B_{0}$ with
@@
 Fix $m\in\mathbb{N}$.
-Let $D\subseteq\mathbb{R}$ be some set with a property that $m^{k}D=m^{k}+D=D$ for any integer $k$.
-Let parameters $(\alpha,\beta)$ be such that $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, and suppose Bob wins $(\alpha,\beta)$-game with $D$.
+Let $S\subseteq\mathbb{R}$ be some set with a property that $m^{k}S=m^{k}+S=S$ for any integer $k$.
+Let parameters $(\alpha,\beta)$ be such that $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, and suppose Bob wins $(\alpha,\beta)$-game with $S$.
 Then for any $\beta^{\prime}<\beta,\,\alpha^{\prime}=\frac{\alpha\beta}{\beta^{\prime}}$, and any closed ball $T$ Bob can win $(\alpha^{\prime},\beta^{\prime})$-game starting with any ball $T$ as his initial ball.
@@
 Now, for $(\alpha^{\prime},\beta^{\prime})$ as in Lemma it remains to take $\varepsilon$ under the condition $(1+\varepsilon)\alpha=\alpha^{\prime}$. ∎
-The discussion above implies the following statement:
+Lemma 3.7 makes the game ”symmetric”: now we can assume that both Alice and Bob can not choose their starting balls. This symmetry makes it possible to utilize symmetric strategies:
 Lemma 3.8.
-Let $m\in\mathbb{N}$ and $D,Q\subseteq\mathbb{R}$ be some sets with the following properties:
-  • $m^{k}D=m^{k}+D=D$ for any integer $k$.
-  • $D\sqcup(-D)\sqcup Q=\mathbb{R}$.
+Let $m\in\mathbb{N}$ and $S,Q\subseteq\mathbb{R}$ be some sets with the following properties:
+  • $m^{k}S=m^{k}+S=S$ for any integer $k$.
+  • $S\sqcup(-S)\sqcup Q=\mathbb{R}$.
 Then:
   1. (1)
-    If $\alpha\geq\beta$, then the set $D$ is $(\alpha,\beta)$-losing.
+    If $\alpha\geq\beta$, then $(\alpha,\beta)$ is a non-winning point for $S$;
   2. (2)
-    If $\beta>\alpha$ and $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, then the set $D\sqcup Q$ is $(\alpha,\beta)$-winning.
+    If $\beta>\alpha$ and $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, then $(\alpha,\beta)$ is a non-losing point for $S\sqcup Q$.
 Proof.
   1. (1)
-    Suppose $\alpha\geq\beta$ and $D$ is $(\alpha,\beta)$-winning.
+    Suppose $\alpha\geq\beta$ and $S$ is $(\alpha,\beta)$-winning.
 In this case Alice has a positional winning strategy described by the function $A(I)$ (where $I\subset\mathbb{R}$ is a closed interval) for this game.
-Consider $(\beta,\alpha)$-game on the set $D$ and let Bob play the following strategy:
+Consider $(\beta,\alpha)$-game on the set $S$ and let Bob play the following strategy:
     • Bob chooses $B_{0}=A([0,1])$.
@@
 Then $-x=\bigcap\limits_{i=0}^{\infty}A(-A_{n})$ is the outcome of the $(\alpha,\beta)$-game where Bob plays the balls $[-1;0];-A_{0},-A_{1},\ldots$ and Alice plays using her winning strategy $A(I)$.
-Therefore, $-x\in D$, which means $-x\notin D$ and Bob wins.
-    It means that the point $(\alpha,\beta)$ is winning and the point $(\beta,\alpha)$ is losing for the set $D$, which contradicts Lemma 3.1. Thus, the point $(\alpha,\beta)$ is losing.
+Therefore, $-x\in S$, which means $-x\notin S$ and Bob wins.
+    It means that the point $(\alpha,\beta)$ is winning and the point $(\beta,\alpha)$ is losing for the set $S$, which contradicts Proposition 3.1. Thus, the point $(\alpha,\beta)$ is non-winning.
   2. (2)
     The idea of the proof is the same as in (a).
-Suppose $D\sqcup Q$ is $(\alpha,\beta)$-losing; Take $\varepsilon>0$ under the condition $(1+\varepsilon)^{2}<\frac{\beta}{\alpha}$.
+Suppose $S\sqcup Q$ is $(\alpha,\beta)$-losing; Take $\varepsilon>0$ under the condition $(1+\varepsilon)^{2}<\frac{\beta}{\alpha}$.
 Then by Lemma 3.7 for any closed ball $T$ Bob has a winning strategy for $\left(\alpha(1+\varepsilon),\frac{\beta}{(1+\varepsilon)}\right)$-game starting from $T$.
@@
 Let $x$ be the outcome of this $(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon))$-game; then $-x$ is the outcome of $(\alpha(1+\varepsilon),\frac{\beta}{(1+\varepsilon)})$-game Bob played with his winning strategy, i.e.
-$-x\in(-D)$ and $x\in D\subseteq D\sqcup Q$.
-It shows that $(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon))$ is a winning point for $D\sqcup Q$.
-    But one has $\alpha<\alpha(1+\varepsilon))<\frac{\beta}{(1+\varepsilon)}$, which by Lemma 3.1 contradicts our initial assumption. Therefore $D\sqcup Q$ is $(\alpha,\beta)$-winning.
-Now, the goal of the next part is to find the condition under which we can remove dyadic rationals from the set $D$ so that it remains winning for the parameters above.
-We start with the following almost obvious statement:
+$-x\in(-S)$ and $x\in S\subseteq S\sqcup Q$.
+It shows that $(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon))$ is a winning point for $S\sqcup Q$.
+    But one has $\alpha<\alpha(1+\varepsilon)<\frac{\beta}{(1+\varepsilon)}$, which by Proposition 3.1 contradicts our initial assumption. Therefore $S\sqcup Q$ is $(\alpha,\beta)$-non-losing.
+Remark 3.9.
+In general, it’s not guaranteed that the game on $S$ as in Lemma 3.8 is determined, so we can only prove the weaker statements about $(\alpha,\beta)$ being non-winning or non-losing.
+Indeed: by ergodicity of the multiplication by $m$ map, one can see that such a $S$ has to have measure zero or be non-measurable.
+In the latter case, such a $S$ is not compatible with the Axiom of determinacy.
+It is not clear if there exists such a non-measurable $S$ on which the game is determined; we only were able to construct the examples of such $S$ using the axiom of choice.
+We finish this section with an almost obvious technical statement that we need to prove Proposition 4.3.
 Lemma 3.10.
-Suppose $S$ is an $(\alpha,\beta)$-winning set and for some $\xi\in\mathbb{R}$ one has $S+\xi\subseteq D$. Then the set $D$ is $(\alpha,\beta)$-winning.
+Suppose $E$ is an $(\alpha,\beta)$-winning set and for some $\xi\in\mathbb{R}$ one has $E+\xi\subseteq S$. Then the set $S$ is $(\alpha,\beta)$-winning.
 Proof.
-If $A(I)$ was Alice’s winning positional strategy on the set $S$, then $A(I-\xi)+\xi$ is a winning positional strategy on the set $D$. ∎
-Now, we use the lemmas we have to prove the following statement:
-Theorem 3.10.
-Fix $m\in\mathbb{Z}$. Let $D\subseteq\mathbb{R}$ be some set with the following properties:
-  • $m^{k}D=m^{k}+D=D$ for any integer $k$.
-  • $D\sqcup(-D)\sqcup Q=\mathbb{R}$, where $Q\subseteq\mathbb{Q}$ is some set.
-  • $D+\xi\subseteq D$ for some $\xi\notin\mathbb{Q}$.
-Then:
-  1. (1)
-    If $\alpha\geq\beta$, then the set $D$ is $(\alpha,\beta)$-losing.
-  2. (2)
-    If $\beta>\alpha$ and $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, then the set $D$ is $(\alpha,\beta)$-winning.
-Proof.
-  1. (1)
-    follows from Lemma 3.8.
-  2. (2)
-    By Lemma 3.8 we know that the set $D\sqcup Q$ is $(\alpha,\beta)$-winning.
-As $Q\subseteq\mathbb{Q}$, one has $D\sqcup Q+\xi\subseteq D$ for any $\xi\notin\mathbb{Q}$.
-By Lemma 3.9 then, $D$ is $(\alpha,\beta)$-winning for $\beta>\alpha$.
-Corollary 3.11.
-Let $D\subseteq\mathbb{R}$ be some set with the following properties:
-  • $m_{i}^{k}D=m_{i}^{k}+D=D$ for any integer $k$ and some $m_{1},\,m_{2}\in\mathbb{N}$ with property $\log_{m_{1}}(m_{2})\notin\mathbb{Q}$;
-  • $D\sqcup(-D)\sqcup Q=\mathbb{R}$, where $Q\subseteq\mathbb{Q}$ is some set;
-  • $D+\xi\subseteq D$ for some $\xi\notin\mathbb{Q}$.
-If $\beta>\alpha$, then the set $D$ is $(\alpha,\beta)$-winning.
-Indeed: for any parameters $(\alpha,\beta)$ in this case either $\log_{m_{1}}(\alpha\beta)$ or $\log_{m_{2}}(\alpha\beta)$ is irrational, so by Theorem 3.10 $D$ is $(\alpha,\beta)$-winning.
-4. Example: sets with a nontrivial diagram
-Theorem 3.10 and Corollary 3.11 give the complete description of the nontrivial Schmidt diagram for sets with special properties.
-The goal of this section is to give an example of a set satisfying these properties, thus proving Theorem 2.1.
-Fix $\xi\notin\mathbb{Q}$.
-Then, we define an auxiliary set $\Xi$ by $\Xi=\{r+q\xi:\,\,r,q\in\mathbb{Q},\,q>0\}$.
-One can see that $-\Xi=\{r+q\xi:\,\,r,q\in\mathbb{Q},\,q<0\}$ is disjoint with $\Xi$, and one has $\Xi+\xi\subseteq\Xi$ by construction.
-Now, let us focus on the set $E=\mathbb{R}\setminus\Big{(}\mathbb{Q}\cup\Xi\cup(-\Xi)\Big{)}$.
-We will divide it into the union of the sets $\overline{x}=\{r+q\xi+px:\,\,r,q,p\in\mathbb{Q},\,\,p\neq 0\}$ over $x\in E$.
-Easy to see that $\overline{x}$ is the equivalence class under relation $\sim$ defined as follows: $x\sim y$ if $y=r+q\xi+px$ with the restrictions as above, therefore $E$ is the disjoint union of the sets $\overline{x}$ over pairwise non-equivalent $x$.
-Next, let $\underline{x}=\{r+q\xi+px:\,\,r,q,p\in\mathbb{Q},\,\,p>0\}$.
-Then $\overline{x}=\underline{x}\sqcup\underline{-x}$.
-Indeed: if $\underline{x}\,\text{and}\,\underline{-x}$ intersect, then $x\in\Xi\sqcup(-\Xi)$, which we do not consider.
-Finally, note that $\underline{-x}=-\underline{x}$.
-Let us enumerate all the sets $\overline{x}$ by some index set $\mathscr{A}$ and denote the corresponding set $E_{\alpha}$ where $\alpha\in\mathscr{A}$.
-Choose some $x_{\alpha}\in A_{\alpha}$ and denote $\underline{x_{\alpha}}=C_{\alpha}$; then $C_{\alpha}\sqcup(-C_{\alpha})=E_{\alpha}$.
-Therefore for $C=\bigsqcup\limits_{\alpha\in\mathscr{A}}C_{\alpha}$ one has $C\sqcup(-C)=E$; by construction also $C+\xi_{j}=C$ (as this works for each of $\underline{x}$).
-Finally, we define $\mathscr{S}=\mathscr{S}(\xi)=C\sqcup\Xi$. Then $-\mathscr{S}=(-C)\sqcup(-\Xi)$, and one has $\mathscr{S}\sqcup(-\mathscr{S})\sqcup\mathbb{Q}=\mathbb{R}$. One also has:
-  • $m^{k}\mathscr{S}=m^{k}+\mathscr{S}=\mathscr{S}$ for any integer $k$ and any $m\in\mathbb{N}$;
-  • $\mathscr{S}+\xi\subseteq\mathscr{S}$.
-Therefore, by Theorem 3.10 and Corollary 3.11, $\mathscr{S}$ is $(\alpha,\beta)$-winning for $\beta>\alpha$ and $(\alpha,\beta)$-losing for $\beta\leq\alpha$. This proves Theorem 2.1.
-5. Digit frequencies
-Next, we will investigate a separate but similar construction regarding frequencies of digits in base-$2$ expansions of real numbers.
-It is known that these sets are closely related to density and dimension, and hence they have been studied in the past (see [V59, FFNS21]).
-First, some setup.
-Consider the base-$2$ expansions of the form $x=x_{0}.x_{1}x_{2}\cdots$ where $x_{0}$ is an integer and $x_{i}\in\{0,1\}$ are the digits in the base-$2$ expansion of $x$.
-In this paper, if $x$ can be represented by two seperate expansions, we take the lexicographically minimal expansion (for example, $0.100\ldots$ would instead be written as $0.011\ldots$).
-We define
-$d^{+}(x,j)=\limsup_{k\rightarrow\infty}\frac{\#\{1\leq i\leq k:x_{i}=j\}}{k}$
-and similarly,
-$d^{-}(x,j)=\liminf_{k\rightarrow\infty}\frac{\#\{1\leq i\leq k:x_{i}=j\}}{k}.$
-Define the sets
-$D^{+}_{c}$ :=$:=\{x\in{\mathbb{R}}:d^{+}(x,0)>c\},$
-$D^{-}_{c}$ :=$:=\{x\in{\mathbb{R}}:d^{-}(x,0)>c\}.$
-The following is known about the Schmidt diagrams of these sets.
-Theorem 5.1.
-[FFNS21] The set $D^{+}_{c}$ is winning for $\log(8\alpha)/\log(\alpha\beta)>c$ and losing for $\log(\alpha/8)/\log(\alpha\beta)\leq c$.
-The results of our research expand on that bound. To do so, we require the following lemma, known as the Equidistribution Theorem:
+If $A(I)$ was Alice’s winning positional strategy on the set $E$, then $A(I-\xi)+\xi$ is a winning positional strategy on the set $S$. ∎
+4. Proof of Theorem 2.2
+We start with the following classical statement about uniform distribution:
 Lemma 4.1.
 [B09] If $\alpha$ is irrational, then the sequence $\alpha,2\alpha,3\alpha,\ldots$ is uniformly distributed modulo $1$.
-Theorem 5.3.
-Suppose $\log_{2}{\alpha\beta}\notin\mathbb{Q}$.
-Then, the sets $D^{\pm}_{c}$ are winning for $\alpha\leq\frac{1}{4}$ and $\log(4\alpha)/\log(\alpha\beta)>c$, and losing for $\beta\leq\frac{1}{4}$ and $\log(\alpha/4)/\log(\alpha\beta)\leq c$.
-In the special case the bounds given by Theorem 5.3 are presented on Figure 5.
+We will prove two separate statements, which together imply the result of Theorem 2.2: Proposition 4.2 and Proposition 4.3.
+Proposition 4.2.
+Let $c\in(0,1)$, and suppose $\log_{2}{\alpha\beta}\notin\mathbb{Q}$.
+Then, the sets $F^{\pm}_{c}$ are $(\alpha,\beta)$-winning if $\log(4\alpha)/\log(\alpha\beta)>c$, and $(\alpha,\beta)$-losing if $\log(\alpha/4)/\log(\alpha\beta)\leq c$.
 Proof.
-We will show the winning strategy for Alice.
-Suppose Bob begins by playing an interval with radius $r_{0}$.
-Alice will play in the following way: suppose Bob plays some interval $B_{t}$.
-Define $k_{t}=-\left\lfloor\log_{2}{(\alpha\beta)^{t}r_{0}}\right\rfloor$, and pick the smallest $m\in\mathbb{Z}$ such that $m/2^{k_{t}}\in B_{t}$.
-Then, it follows that the dyadic interval $\left(\frac{m}{2^{k_{t}}},\frac{m+1}{2^{k_{t}}}\right)\subset B_{t}$.
-Then, Alice plays $A_{t}=\left(\frac{m}{2^{k_{t}}},\frac{m}{2^{k_{t}}}+2\alpha(\alpha\beta)^{t}r_{0}\right)$.
+First, let us notice that our winning conditions imply that $\alpha<\frac{1}{4}$, and the losing conditions imply that $\beta\leq\frac{1}{4}$.
+We will show the winning strategy under these conditions for Alice in the $(\alpha,\beta)$-game on the set $F^{-}_{c}$ (and thus on $F^{+}_{c}$, since $F^{-}_{c}\subseteq F^{+}_{c}$).
+Let $r_{0}$ be the radius of $B_{0}:\,\,r_{0}=\frac{1}{2}|B_{0}|$, and suppose Bob played some interval $B_{t}$.
+Define $k_{t}=-\left\lfloor\log_{2}{(\alpha\beta)^{t}r_{0}}\right\rfloor$.
+Consider the smallest $m\in\mathbb{Z}$ such that $m/2^{k_{t}}\in B_{t}$.
+For such an $m$, one has $\left[\frac{m}{2^{k_{t}}},\frac{m+1}{2^{k_{t}}}\right]\subset B_{t}$.
+Then, Alice plays $A_{t}=\left[\frac{m}{2^{k_{t}}},\frac{m}{2^{k_{t}}}+2\alpha(\alpha\beta)^{t}r_{0}\right]$.
 Define $l_{t}$ such that $k_{t}+l_{t}=-\left\lfloor\log_{2}{\alpha(\alpha\beta)^{t}r_{0}}\right\rfloor-2$.
-Then, Alice ensures that for any $x\in A_{t}$, the $(k_{t}+1)$th digit to the $(k_{t}+l_{t})$th digit is equal to 0.
+Then, Alice ensured by her move that for any $x\in A_{t}$, the $(k_{t}+1)$-th to the $(k_{t}+l_{t})$-th digits are equal to 0.
 In other words, she guarantees $l_{t}$ zeroes on this turn.
-For convenience, let $a=\log_{2}\alpha$, $b=\log_{2}\beta$, and $\rho=\log_{2}r_{0}$; note that $a+b\notin\mathbb{Q}$.
+For convenience, let $a=\log_{2}\alpha$, $b=\log_{2}\beta$, and $\rho=\log_{2}r_{0}$; note that $a+b\notin\mathbb{Q}$, and $a+2<0$.
 We have
@@
 $=-(a+2)+\{a+(a+b)t+\rho\}-\{(a+b)t+\rho\}.$
-This implies that by the Equidistribution Theorem,
+This implies that by the Lemma 4.1,
 $\lim\limits_{n\rightarrow\infty}\frac{\sum_{i=0}^{n-1}l_{i}}{n}$ $=-(a+2).$
-Now, notice that for $k_{n}\leq k<k_{n+1}$ one has
+Now, let $x=x_{0}+\sum\limits_{i=1}^{\infty}x_{i}2^{-i}=\bigcap\limits_{t=0}^{\infty}A_{t}$ be the outcome of the game. We notice that for $k_{n}\leq k<k_{n+1}$ one has
 #$\frac{\#\{1\leq i\leq k:x_{i}=0\}}{k}=\frac{\#\{1\leq i\leq k_{n}:x_{i}=0\}}{k_{n}}\cdot\frac{k_{n}}{k}+\frac{\#\{k_{n}+1\leq i\leq k:x_{i}=0\}}{k},$
@@
 $=\frac{-(a+2)}{-(a+b)}$
-$=\frac{\log(4\alpha)}{\log(\alpha\beta)}>c$
-as desired. Therefore, Alice wins. The losing condition is analogous, so we omit it.
-For the proofs below, we will use the following
+$=\frac{\log(4\alpha)}{\log(\alpha\beta)}>c,$
+which means that $x\in F^{-}_{c}$. Therefore, Alice wins. The losing condition is analogous, so we omit it.
+For the next statement, we fix $c=\frac{1}{2}$.
+Together with the proposition above, it clearly implies the statement of Theorem 2.2 due to trivial inclusions: $F_{c_{1}}^{\pm}\subseteq F_{c_{2}}^{\pm}$ if $c_{1}\geq c_{2}$.
+Proposition 4.3.
+  1. (1)
+    If $\alpha\geq\beta$, then the set $F^{-}_{\frac{1}{2}}=\{x:\,d^{-}(x,0)>\frac{1}{2}\}$ is $(\alpha,\beta)$-losing;
+  2. (2)
+    If $\beta>\alpha$ and $\log_{2}(\alpha\beta)\notin\mathbb{Q}$, then the set $\{x:\,d^{+}(x,0)\geq\frac{1}{2}\}$ is $(\alpha,\beta)$-winning.
+In order to show Proposition 4.3, we will need the following observation:
 Lemma 4.4.
-Let $S\subseteq\mathbb{N}$ be an infinite set of density zero. Let $d^{\pm}(x,0)=c$; then
-$d^{\pm}\left(x+\sum\limits_{j\in S}\frac{1}{2^{j}},0\right)\geq c.$
+Let $\mathfrak{S}\subseteq\mathbb{N}$ be an infinite set of density zero, and define $\xi:=\sum\limits_{j\in\mathfrak{S}}\frac{1}{2^{j}}$. Suppose $d^{\pm}(x,0)=c$; then
+$d^{\pm}\left(x+\xi,0\right)\geq c.$
 Proof.
-For $x\in\mathbb{Z}[\frac{1}{2}]$ one has $d^{\pm}t(x+\sum\limits_{j\in S}\frac{1}{2^{j}},0)=d^{\pm}(\sum\limits_{j\in S}\frac{1}{2^{j}},0)=1$, as addition of $x$ changes only finitely many digits.
+For $x\in\mathbb{Z}[\frac{1}{2}]$ one has $d^{\pm}t(x+\sum\limits_{j\in\mathfrak{S}}\frac{1}{2^{j}},0)=d^{\pm}(\sum\limits_{j\in\mathfrak{S}}\frac{1}{2^{j}},0)=1$, as addition of $x$ changes only finitely many digits.
 Now assume $x\notin\mathbb{Z}[\frac{1}{2}]$.
 […]
-$d^{\pm}(x+\sum\limits_{j\in S_{2}}\frac{1}{2^{j}},0)={\lim\limits_{k\rightarrow\infty}}^{\pm}\frac{\#\{1\leq i\leq k:\,i\notin Q\sqcup S_{2}\}}{k}=$
-${\lim\limits_{k\rightarrow\infty}}^{\pm}\Big{(}\frac{\#\{1\leq i\leq k:\,i\notin Q\}}{k}-\frac{\#\{1\leq i\leq k:\,i\in S_{2}\}}{k}\Big{)}=$
-${\lim\limits_{k\rightarrow\infty}}^{\pm}\frac{\#\{1\leq i\leq k:\,i\notin Q\}}{k}-\lim\limits_{k\rightarrow\infty}\frac{\#\{1\leq i\leq k:\,i\in S_{2}\}}{k}=$
+$d^{\pm}(x+\sum\limits_{j\in\mathfrak{S}_{2}}\frac{1}{2^{j}},0)={\lim\limits_{k\rightarrow\infty}}^{\pm}\frac{\#\{1\leq i\leq k:\,i\notin Q\sqcup\mathfrak{S}_{2}\}}{k}=$
+${\lim\limits_{k\rightarrow\infty}}^{\pm}\Big{(}\frac{\#\{1\leq i\leq k:\,i\notin Q\}}{k}-\frac{\#\{1\leq i\leq k:\,i\in\mathfrak{S}_{2}\}}{k}\Big{)}=$
+${\lim\limits_{k\rightarrow\infty}}^{\pm}\frac{\#\{1\leq i\leq k:\,i\notin Q\}}{k}-\lim\limits_{k\rightarrow\infty}\frac{\#\{1\leq i\leq k:\,i\in\mathfrak{S}_{2}\}}{k}=$
 $={\lim\limits_{k\rightarrow\infty}}^{\pm}\frac{\#\{1\leq i\leq k:\,i\notin Q\}}{k}-0=d^{\pm}(x,0)=c.$
-Now it remains to prove the statement only for $S$ such that $S\subseteq Q$.
-Let us define $R\subseteq\mathbb{N}$ by $\sum\limits_{j\in Q}\frac{1}{2^{j}}+\sum\limits_{j\in S}\frac{1}{2^{j}}=\sum\limits_{j\in R}\frac{1}{2^{j}}$.
+Now it remains to prove the statement only for $\mathfrak{S}$ such that $\mathfrak{S}\subseteq Q$.
+Let us define $R\subseteq\mathbb{N}$ by $\sum\limits_{j\in Q}\frac{1}{2^{j}}+\sum\limits_{j\in\mathfrak{S}}\frac{1}{2^{j}}=\sum\limits_{j\in R}\frac{1}{2^{j}}$.
 It is easy to check that:
@@
   • Suppose $a\notin Q,\,a+b\notin Q$ and $i\in Q$ for all $a<i<a+b$. Then #$\#\{a\leq i<a+b:\,i\in R\}\leq b-1=\#\{a\leq i<a+b:\,i\in Q\}$.
-This immediately implies that $d^{\pm}(x+\sum\limits_{j\in S}\frac{1}{2^{j}},0)\geq d^{\pm}(x,0)=c$ and completes the proof. ∎
+This immediately implies that $d^{\pm}(x+\sum\limits_{j\in\mathfrak{S}}\frac{1}{2^{j}},0)\geq d^{\pm}(x,0)=c$ and completes the proof. ∎
 Let us also recall simple properties of digit frequencies:
+  1. (1)
     $d^{\pm}(2^{k}x,i)=d^{\pm}(2^{l}+x,i)$ for any integers $k,l$;
+  2. (2)
     $d^{\pm}(x,i)+d^{\mp}(x,1-i)=1$;
+  3. (3)
     If $x\notin\mathbb{Z}[\frac{1}{2}]$, then $d^{\pm}(x,i)=d^{\pm}(-x,1-i)$.
-We now focus on the special case $c=\frac{1}{2}$. We will use the construction similar to the set $\mathscr{D}$ but connected to digit frequencies.
-Let $\mathscr{F}=D^{-}_{\frac{1}{2}}\cup(\mathscr{D}(\xi)\cap\{x:\,d^{+}(x,0)\geq\frac{1}{2}\})$ where $\xi=\sum\limits_{i=1}^{\infty}\frac{1}{2^{2^{i}}}$.
-We want to show that the set $\mathscr{F}$ satisfies the conditions of Theorem 3.10 for $m=2$.
-Proposition 5.5.
+We are ready to prove Proposition 4.3.
+Proof of Proposition 4.3.
+Let $S=F_{\frac{1}{2}^{-}}\setminus\mathbb{Z}[\frac{1}{2}]$.
+Then, using the properties 2 and 3 listed above, it is easy to verify that $-S=\{x:\,\,d^{+}(-x,0)<\frac{1}{2}\}\setminus\mathbb{Z}[\frac{1}{2}]$ is a set disjoint with $S$.
+Let us define $Q$ under the condition $S\sqcup(-S)\sqcup Q=\mathbb{R}$; in this case, $S\sqcup Q=\{x:\,d^{+}(x,0)\geq\frac{1}{2}\}\cup\mathbb{Z}[\frac{1}{2}]$.
+Property 1 shows that the sets $S$ and $Q$ satisfy the conditions of the Lemma 3.8 with $m=2$. As in addition our sets are Borel, the $(\alpha,\beta)$-game on them is determined, and we can claim that
+  • If $\alpha\geq\beta$, then $(\alpha,\beta)$ is a losing point for $S$;
+  • If $\beta>\alpha$ and $\log_{2}(\alpha\beta)\notin\mathbb{Q}$, then $(\alpha,\beta)$ is a winning point for $S\sqcup Q$.
+Now let us fix $\xi$ as in Lemma 4.4: $\xi$ is irrational, so by the aforementioned lemma $F_{\frac{1}{2}}^{-}+\xi\subseteq S$ and $S\sqcup Q+\xi\subseteq\{x:\,d^{+}(x,0)\geq\frac{1}{2}\}$.
+The following two observations complete the proof of Proposition 4.3:
   1. (1)
-    $2^{k}\mathscr{F}=2^{k}+\mathscr{F}=\mathscr{F}$ for any integer $k$;
+    Suppose $F_{\frac{1}{2}}^{-}$ is $(\alpha,\beta)$-winning for $\alpha\geq\beta$; then, by Lemma 3.10, $S$ should also be $(\alpha,\beta)$-winning, which is not true.
+Provided that the game is determined on $F_{\frac{1}{2}}^{-}$, we conclude that the set $F_{\frac{1}{2}}^{-}$ is $(\alpha,\beta)$-losing.
   2. (2)
-    $\mathscr{F}\sqcup(-\mathscr{F})\sqcup Q=\mathbb{R}$ where $Q\subseteq\mathbb{Q}$;
-  3. (3)
-    $\mathscr{F}+\xi\subseteq\mathscr{F}$.
-Proof.
-  1. (1)
-    Follows from the similar properties of all the sets in construction of $\mathscr{F}$;
-  2. (2)
-    It’s easy to notice that
-    =$-\mathscr{F}=\left\{x\notin\mathbb{Z}\left[\frac{1}{2}\right]:\,d^{+}(x,0)<\frac{1}{2}\right\}\cup\left((-\mathscr{D}(\xi))\cap\left\{x:\,d^{-}(x,0)\leq\frac{1}{2}\right\}\right),$
-    which shows that $\mathscr{F}$ and $-\mathscr{F}$ are disjoint.
-If $x\notin\mathscr{F}\sqcup(-\mathscr{F})$, then either $x\in\mathbb{Z}[\frac{1}{2}]$ or $d^{-}(x,0)\leq\frac{1}{2}\leq d^{+}(x,0)$ and $x\notin\mathscr{D}\sqcup(-\mathscr{D})$, which implies $x\in\mathbb{Q}$.
-  3. (3)
-    By Lemma 5.4 all the sets in the construction of $\mathscr{F}$ satisfy this condition.
-This immediately proves that
-  • If $\alpha\geq\beta$, then the set $\mathscr{F}$ is $(\alpha,\beta)$-losing;
-  • If $\beta>\alpha$ and $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, then the set $\mathscr{F}$ is $(\alpha,\beta)$-winning.
-Let’s also note that $\{x:\,d^{-}(x,0)>\frac{1}{2}\}\subseteq\mathscr{F}\subseteq\{x:\,d^{+}(x,0)\geq\frac{1}{2}\}$, thus
-Corollary 5.6.
-  1. (1)
-    If $\alpha\geq\beta$, then the set $D^{-}_{\frac{1}{2}}=\{x:\,d^{-}(x,0)>\frac{1}{2}\}$ is $(\alpha,\beta)$-losing;
-  2. (2)
-    If $\beta>\alpha$ and $\log_{2}(\alpha\beta)\notin\mathbb{Q}$, then the set $\{x:\,d^{+}(x,0)\geq\frac{1}{2}\}$ is $(\alpha,\beta)$-winning.
-Conjecture 5.7.
-We believe that $D(D^{-}_{c})=\{(\alpha,\beta):\,\log(\alpha)/\log(\alpha\beta)>c\}$.
-6. B-Badly Approximable Numbers
-The set of badly approximable numbers is an important object of study in Diophantine approximations. It is defined as follows:
-$\operatorname{BA}:=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{n}\right|>\frac{c}{n^{2}}\text{ for some $c>0$ and all }m\in{\mathbb{Z}},\ n\in{\mathbb{N}}\right\}.$
-It is known that the badly approximable numbers are dense and have measure zero.
-Interestingly enough, despite having measure zero, Schmidt showed that they are winning for the largest possible nontrivial bounds:
-Theorem 6.1.
-[S66] The set $\operatorname{BA}$ is winning for all $(\alpha,\beta)\in\hat{D}$; that is, $D(\operatorname{BA})=\hat{D}$.
-In a similar vein to the badly approximable numbers investigated by Schmidt, we define the set of B-Badly Approximable Numbers as
-$\operatorname{B-BA}:=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{B^{k}}\right|>\frac{c}{B^{k}}\text{ for some $c>0$ and all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\right\}$
-and a closely related set
-$\operatorname{B-BA}(c):=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{B^{k}}\right|>\frac{c}{B^{k}}\text{ for all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\right\}.$
-These definitions imply $\operatorname{B-BA}=\bigcup_{c>0}\operatorname{B-BA}(c)$. We note that the $\operatorname{B-BA}$ set has been previously studied in [N10] and [PS08].
-The following statement follows from work of Schmidt:
-Theorem 6.2.
-[S66] The set $\operatorname{2-BA}$ is winning for all $(\alpha,\beta)\in\hat{D}$; that is, $D(\operatorname{2-BA})=\hat{D}$.
-While both $\operatorname{2-BA}$ and $\operatorname{2-BA}(c)$ did not produce any new Schmidt diagrams, there is a simple change we can make that will produce both nontrivial winning and nontrivial losing zones.
-For any integer $B\geq 2$, define
-$\operatorname{\widetilde{B-BA}}(c,N):=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{B^{k}}\right|>\frac{c}{B^{k}}\text{ for all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\ \text{s.t.}\ k>N\right\}$
-and
-$\operatorname{\widetilde{B-BA}}(c):=\bigcup_{N\in{\mathbb{N}}}\operatorname{\widetilde{B-BA}}(c,N).$
-Throughout this chapter, we define $\gamma_{A}=1-2\alpha+\alpha\beta$ and $\gamma_{B}=1-2\beta+\alpha\beta$. We begin results in this section with a useful general lemma:
+    Follows directly from Lemma 3.10 and observations above.
+5. Proofs of the statements regarding d-Badly Approximable Numbers
+We begin results in this section with a useful general lemma:
 Lemma 5.1.
-Suppose Bob plays any interval $B_{t}$ and $\gamma_{A}>0$.
+Suppose Bob plays any interval $B_{t}$ and $\gamma>0$.
 Denote the center of $B_{t}$ by $c_{t}$.
-Then, Bob can always choose an interval $B_{t+1}$ such that $|c_{t+1}-c_{t}|\leq\frac{|B_{t}|}{2}\gamma_{A}$.
-As a corollary, Bob can guarantee that for any $m\in\mathbb{N}$, $|c_{t+m}-c_{t}|\leq\frac{|B_{t}|}{2(1-\alpha\beta)}\gamma_{A}$.
-Analogous results hold for Alice, with $\gamma_{A}$ replaced by $\gamma_{B}$.
+Then, Bob can always choose an interval $B_{t+1}$ such that $|c_{t+1}-c_{t}|\leq\frac{|B_{t}|}{2}\gamma$.
+As a corollary, Bob can guarantee that for any $m\in\mathbb{N}$, $|c_{t+m}-c_{t}|\leq\frac{|B_{t}|}{2(1-\alpha\beta)}\gamma$.
+Analogous results hold for Alice, with $\gamma$ replaced by $1-2\beta+\alpha\beta$.
 Proof.
@@
 Let’s denote center of $B_{t}$ by $c_{t}$.
-It is clear that Bob can always choose such a $B_{t+1}$ that $|c_{r+1}-c_{r}|\leq\frac{|B_{t}|}{2}\gamma_{A}$, since no matter which side Alice picks her interval on, Bob will go the opposite way.
-Extending this logic, it follows that $|c_{m+r}-(a_{0}+k\delta)|\leq\frac{\rho(\alpha\beta)^{m}}{2}\gamma_{A}(1+(\alpha\beta)+(\alpha\beta)^{2}+\ldots)=\frac{\rho(\alpha\beta)^{m}}{2(1-\alpha\beta)}\gamma_{A}$, which completes the proof.
-The following theorems provide the winning and losing conditions for $\operatorname{\widetilde{B-BA}}$ sets.
-Theorem 6.4.
-If
-$\frac{2c(1-\beta)(1-\alpha\beta)}{\beta\gamma_{A}}>1,$
-then for any $B$, the set $\operatorname{\widetilde{B-BA}}(c)$ is $(\alpha,\beta)$-losing.
-Proof.
+It is clear that Bob can always choose such a $B_{t+1}$ that $|c_{r+1}-c_{r}|\leq\frac{|B_{t}|}{2}\gamma$, since no matter which side Alice picks her interval on, Bob will go the opposite way.
+Extending this logic, it follows that $|c_{m+r}-(a_{0}+k\delta)|\leq\frac{\rho(\alpha\beta)^{m}}{2}\gamma(1+(\alpha\beta)+(\alpha\beta)^{2}+\ldots)=\frac{\rho(\alpha\beta)^{m}}{2(1-\alpha\beta)}\gamma$, which completes the proof.
+Now we are ready to give the proofs of our main theorems.
+Proof of Theorem 2.5
 One has
-$\operatorname{\widetilde{B-BA}}(c)^{c}=\left\{x:\,\,\forall N\,\,\exists\,n>N\,\,\text{and}\,\,k\,\in\,\mathbb{Z}:\,\,\left|x-\frac{k}{B^{n}}\right|<\frac{c}{B^{n}}\right\}=$
-$=\bigcap\limits_{N=1}^{\infty}\bigcup\limits_{n=N}^{\infty}\left(\bigcup\limits_{k\in\mathbb{Z}}\left[\frac{k}{B^{n}}-\frac{c}{B^{n}},\frac{k}{B^{n}}+\frac{c}{B^{n}}\right]\right)\subseteq\bigcap\limits_{N=1}^{\infty}\bigcup\limits_{n=N}^{\infty}\left(\bigcup\limits_{k\in\mathbb{Z}}\left(\frac{k}{B^{n}}-\frac{c}{B^{n}},\frac{k}{B^{n}}+\frac{c}{B^{n}}\right)\right),$
-and let $J_{n}=\bigcup\limits_{k\in\mathbb{Z}}\left(\frac{k}{B^{n}}-\frac{c}{B^{n}},\frac{k}{B^{n}}+\frac{c}{B^{n}}\right)$. Fix $\varepsilon>0$ such that
-$\varepsilon=\frac{2c(1-\beta)(1-\alpha\beta)}{\beta\gamma_{A}}-1.$
+$\operatorname{\widetilde{d-BA}}(c)^{c}=\left\{x:\,\,\forall N\,\,\exists\,n>N\,\,\text{and}\,\,k\,\in\,\mathbb{Z}:\,\,\left|x-\frac{k}{d^{n}}\right|\leq\frac{c}{d^{n}}\right\}=$
+$=\bigcap\limits_{N=1}^{\infty}\bigcup\limits_{n=N}^{\infty}\left(\bigcup\limits_{k\in\mathbb{Z}}\left[\frac{k}{d^{n}}-\frac{c}{d^{n}},\frac{k}{d^{n}}+\frac{c}{d^{n}}\right]\right)\subseteq\bigcap\limits_{N=1}^{\infty}\bigcup\limits_{n=N}^{\infty}\left(\bigcup\limits_{k\in\mathbb{Z}}\left(\frac{k}{d^{n}}-\frac{c}{d^{n}},\frac{k}{d^{n}}+\frac{c}{d^{n}}\right)\right),$
+and let $J_{n}=\bigcup\limits_{k\in\mathbb{Z}}\left(\frac{k}{d^{n}}-\frac{c}{d^{n}},\frac{k}{d^{n}}+\frac{c}{d^{n}}\right)$. Fix $\varepsilon>0$ such that
+$\varepsilon=\frac{2c(1-\beta)(1-\alpha\beta)}{\beta\gamma}-1.$
 Bob will play in the following way:
   • On his first turn, Bob will play an interval of length
-    $\frac{1}{B}\left(\frac{2c(1-\alpha\beta)}{\gamma_{A}}\right)$
-    centered at $1/B$. By Lemma 1.2, Bob can play in such a way that in finitely many moves, $B_{s}\subset J_{1}$.
+    $\frac{1}{d}\left(\frac{2c(1-\alpha\beta)}{\gamma}\right)$
+    centered at $1/d$. By Lemma 1.2, Bob can play in such a way that in finitely many moves, $B_{s}\subset J_{1}$.
   • On all subsequent turns, Bob plays arbitrarily until some turn $t$ such that
-    $\log_{B}\left(\frac{1}{1-\beta}\right)$$\leq\left\{\left(t\cdot\log_{B}\alpha\beta\right)+\log_{B}|A_{0}|\right\}$ (6.1)
-    $\leq\log_{B}\left(\frac{1}{1-\beta}\right)+\log_{B}(1+\varepsilon),$
+    $\log_{d}\left(\frac{1}{1-\beta}\right)$$\leq\left\{\left(t\cdot\log_{d}\alpha\beta\right)+\log_{d}|A_{0}|\right\}$ (5.1)
+    $\leq\log_{d}\left(\frac{1}{1-\beta}\right)+\log_{d}(1+\varepsilon),$
     there exists $N\in\mathbb{N}$ such that
-    $\left(\frac{1}{1-\beta}\right)\leq B^{N}|A_{t}|\leq\left(\frac{1}{1-\beta}\right)(1+\varepsilon)=\left(\frac{2c(1-\alpha\beta)}{\beta\gamma_{A}}\right),$ (6.2)
-    and since $|A_{t}|(1-\beta)>1/B^{N}$, Bob can play his next interval centered at $m/B^{N}$ for some $m\in\mathbb{Z}$.
+    $\left(\frac{1}{1-\beta}\right)\leq d^{N}|A_{t}|\leq\left(\frac{1}{1-\beta}\right)(1+\varepsilon)=\left(\frac{2c(1-\alpha\beta)}{\beta\gamma}\right),$ (5.2)
+    and since $|A_{t}|(1-\beta)>1/d^{N}$, Bob can play his next interval centered at $m/d^{N}$ for some $m\in\mathbb{Z}$.
 Furthermore, by Lemma 1.2, Bob can play in such a way that in finitely many moves, $B_{t+s}\subset J_{n}$.
-By the Equidistribution Theorem, there exists an infinite number of possible $t$ that satisfy inequality 6.2.
+By Lemma 4.1, there exists an infinite number of possible $t$ that satisfy inequality 5.2.
 Therefore, Bob can do this an infinite number of times, and Alice loses.
-In the special case of the $\operatorname{\widetilde{2-BA}}$ set, we have the following theorem.
-Theorem 6.5.
-The set $\operatorname{\widetilde{2-BA}}(c)$ is winning for $\alpha<\frac{1}{3}$, $\beta>\frac{9c}{1-3\alpha}$, and $c<1/6$.
-Proof.
-The winning strategy for Alice is as follows: suppose Alice is given an interval $B_{i}$ of length $l_{i}$. Pick $k_{i}$ such that
-$\frac{1}{2^{k_{i}}}<l_{i}\leq\frac{1}{2^{k_{i}-1}}$
-or equivalently,
-$\frac{l_{i}}{2}\leq\frac{1}{2^{k_{i}}}<l_{i}.$
-Let $W_{i}$ denote the set of all intervals of length $\alpha l_{i}$ and center $c_{i}$ of the form $(m\pm\frac{1}{3})2^{-k_{i}}$, where $m\in{\mathbb{Z}}$.
-Then, Alice will select an interval from $W_{i}$.
-This is always possible because the maximum distance between the two endpoints of sets is
-$\left|c_{i}-c^{\prime}_{i}\right|+\alpha l_{i}=\frac{1}{3\cdot 2^{k_{i}-1}}+\alpha l_{i}<\frac{2l_{i}}{3}+\frac{l_{i}}{3}=l_{i}$
-and thus any interval of length $l_{i}$ must contain at least one of the intervals in $W_{i}$.
-Next, define $q_{i}$ to be the smallest integer such that $m/2^{q_{i}}\in A_{i}$ for some $m\in{\mathbb{Z}}$. Note that
-$\left|\frac{m}{2^{q_{i}}}-c_{i}\right|=\frac{1}{3\cdot 2^{q_{i}}}<\frac{\alpha l_{i}}{2}<\frac{1}{3\cdot 2^{q_{i}-1}}.$
+Proof of Theorem 2.6 We will describe the winning strategy for Alice. Suppose on the step $t$ of the game Bob played an interval $B_{t}$ of length $l_{t}$. Pick $k_{t}$ such that
+$\frac{l_{t}}{2}\leq\frac{1}{2^{k_{t}}}<l_{t}\leq\frac{1}{2^{k_{t}-1}}.$ (5.3)
+Alice’s strategy is to choose a ball of length $\alpha l_{t}$ and center $c_{t}$ of the form $(r\pm\frac{1}{3})2^{-k_{t}}$ for some $r\in{\mathbb{Z}}$ as $A_{t}$.
+She can always do this, as for any two subsequent points $c_{t}^{1},c_{t}^{2}$ of the chosen form one has
+$\left|c^{1}_{t}-c^{2}_{t}\right|+\alpha l_{t}=\frac{1}{3\cdot 2^{k_{t}-1}}+\alpha l_{t}<\frac{2l_{t}}{3}+\frac{l_{t}}{3}=l_{t}.$
+In some special cases, when needed, we will specify which of such balls Alice will need to pick.
+Define $q_{t}$ to be the smallest integer such that $m/2^{q_{t}}\in A_{t}$ for some $m\in{\mathbb{Z}}$.
+Note that $q_{t}>k_{t}$, as no points of the form $\frac{s}{2^{k_{t}}}$ with $s\in\mathbb{Z}$ belong to $A_{t}$ by construction.
+One has
+$\left|\frac{m}{2^{q_{t}}}-c_{t}\right|=\frac{1}{3\cdot 2^{q_{t}}}<\frac{\alpha l_{t}}{2}<\frac{1}{3\cdot 2^{q_{t}-1}}.$ (5.4)
 We wish to show that for all $i\geq 1$ and $x\in A_{i}$,
-$\left|x-\frac{m}{2^{n}}\right|>\frac{c}{2^{n}}\ \text{for all}\ n\leq q_{i}-2.$
-by induction on $i$.
-Suppose that the statement holds for $i$.
-Then, Bob picks an interval $B_{i+1}$ with length $l_{i+1}=\alpha\beta l_{i}$.
-Now, note that for all $x\in A_{i+1}$ and $q_{i}\leq n\leq k_{i+1}$,
+$\left|x-\frac{m}{2^{n}}\right|>\frac{c}{2^{n}}\ \text{for all}\ n\leq q_{i}-2$
+by induction on the number of step $i$. Suppose that the statement holds for $i$, and Bob picks an interval $B_{i+1}$ with length $l_{i+1}=\alpha\beta l_{i}$.
+  • First, we consider $q_{i}\leq n\leq k_{i+1}$. Note that for all $x\in A_{i+1}$ and any $m\in\mathbb{Z}$
     $\left|x-\frac{m}{2^{n}}\right|\geq\left|x-\frac{m^{\prime}}{2^{k_{i+1}}}\right|$ $>\frac{1}{3\cdot 2^{k_{i+1}}}-\frac{\alpha l_{i+1}}{2}$
@@
     $=\frac{3}{2}\alpha l_{i}c$
-$>\frac{c}{2^{q_{i}}}\geq\frac{c}{2^{n}}.$
-To take care of $k_{i+1}+1\leq n\leq q_{i+1}-2$, simply note that for all $x\in A_{i+1}$ and $c<1/6$,
-$\left|x-\frac{m}{2^{n}}\right|$ $>\frac{1}{3\cdot 2^{n}}-\frac{\alpha l_{i+1}}{2}>\frac{1}{3\cdot 2^{n}}-\frac{1}{3\cdot 2^{q_{i+1}-1}}>\frac{c}{2^{n}}.$
-Finally, for $n=q_{i}-1$, we split into cases based on $A_{i}$, the prior move.
+    $>\frac{c}{2^{q_{i}}}\geq\frac{c}{2^{n}},$
+    where we used 5.4 in the second to last inequality.
+  • To take care of $k_{i+1}+1\leq n\leq q_{i+1}-2$, simply note that for any $x\in A_{i+1}$ and any $c<1/6$,
+    $\left|x-\frac{m}{2^{n}}\right|$ $>\frac{1}{3\cdot 2^{n}}-\frac{\alpha l_{i+1}}{2}>\frac{1}{3\cdot 2^{n}}-\frac{1}{3\cdot 2^{q_{i+1}-1}}>\frac{1}{6\cdot 2^{n}}\geq\frac{c}{2^{n}}.$
+    Here, we again used 5.4 in the second inequality.
+  • Finally, consider $n=q_{i}-1$. There are two possible cases based on $A_{i}$, the prior move.
     • Case I.
@@
       $\frac{\alpha l_{i}}{2}<\frac{1}{3\cdot 2^{q_{i}-1}}-\frac{c}{2^{q_{i}-1}},$
-    then for all $x\in A_{i}$,
-    $\left|x-\frac{m}{2^{q_{i}-1}}\right|$ $\geq\frac{1}{3\cdot 2^{q_{i}-1}}-\frac{\alpha l_{i}}{2}$
-    $>\frac{1}{3\cdot 2^{q_{i}-1}}-\frac{1}{3\cdot 2^{q_{i}-1}}+\frac{c}{2^{q_{i}-1}}$
-    $=\frac{c}{2^{q_{i}-1}}.$
+      then for all $x\in A_{i}$ and $m\in\mathbb{Z}$
+      $\left|x-\frac{m}{2^{q_{i}-1}}\right|$ $\geq\frac{1}{3\cdot 2^{q_{i}-1}}-\frac{\alpha l_{i}}{2}>\frac{c}{2^{q_{i}-1}},$
+      which means that the desired inequality is already satisfied for any $x\in A_{i}$ and thus for any $x\in A_{i+1}$.
     • Case II.
@@
       $\frac{\alpha l_{i}}{2}\geq\frac{1}{3\cdot 2^{q_{i}-1}}-\frac{c}{2^{q_{i}-1}}.$
-    Then, if
-    $B_{i+1}\cap\left[\frac{m-c}{2^{q_{i}-1}},\frac{m+c}{2^{q_{i}-1}}\right]=\emptyset,$
-    it is clear that Alice does not have to worry about $n=q_{i}-1$, since for all $x\in A_{i+1}\subset B_{i+1}$, $|x-m/2^{q_{i}-1}|>c/2^{q_{i}+1}$.
-Thus, for the remainder of this case, assume $B_{i+1}$ intersects the interval.
-Without loss of generality, assume that $m/2^{q_{i}-1}>x$ for all $x\in B_{i+1}$ (that is, the fraction lies to the right of the interval $B_{i+1}$).
-This is possible because the mapping $x\mapsto-x$ gives us the other case.
-Alice will then choose an interval $A_{i+1}\in W_{i+1}$ centered at $c_{i+1}$ such that
+      If
+      $B_{i+1}\cap\left[\frac{m-c}{2^{q_{i}-1}},\frac{m+c}{2^{q_{i}-1}}\right]=\emptyset$
+      for any $m\in\mathbb{Z}$, it is clear that Alice does not have to worry about $n=q_{i}-1$, since for all $x\in A_{i+1}\subset B_{i+1}$, $|x-m/2^{q_{i}-1}|>c/2^{q_{i}+1}$.
+Thus, for the remainder of this case, assume $B_{i+1}$ intersects such an interval.
+Without loss of generality (the other case is symmetric), assume also that $m/2^{q_{i}-1}>x$ for all $x\in B_{i+1}$ (that is, the fraction lies to the right of the interval $B_{i+1}$).
+Alice will then choose an interval $A_{i+1}$ according to her strategy which is centered at $c_{i+1}$ such that
       $c_{i+1}\leq\frac{m}{2^{q_{i}-1}}-\frac{2}{3}\left(\frac{1}{2^{k_{i+1}}}\right),$
       that is, Alice can not choose the closest interval to $\frac{m}{2^{q_{i}-1}}$.
-This is always possible because the second closest interval is within $l_{i+1}$ of $\frac{m}{2^{q_{i}-1}}$; that is, for $x\in A_{i+1}$,
-    $\left|x-\frac{m}{2^{q_{i}-1}}\right|<\frac{2}{3}\left(\frac{1}{2^{k_{i+1}}}\right)+\frac{\alpha l_{i+1}}{2}<\frac{2}{3}l_{i+1}+\frac{1}{6}l_{i+1}<l_{i+1}$
-    and thus whenever Alice is able to pick the closest interval, Alice is also able to pick the second closest interval. Furthermore, for $x\in A_{i+1}$,
+This is always possible because the second closest interval of the desired form is within $l_{i+1}$ of $\frac{m}{2^{q_{i}-1}}$.
+Indeed: the left endpoint of the second closest interval of this form is
+      $e_{i+1}=\frac{m}{2^{q_{i}-1}}-\frac{2}{3}\left(\frac{1}{2^{k_{i+1}}}\right)-\frac{\alpha l_{i+1}}{2},$
+      while the left endpoint of $B_{i+1}$ is smaller than $\frac{m}{2^{q_{i}-1}}-l_{i+1}<e_{i}$, where the latter inequality holds as
+      $\frac{2}{3}\left(\frac{1}{2^{k_{i+1}}}\right)+\frac{\alpha l_{i+1}}{2}<\frac{2}{3}l_{i+1}+\frac{1}{6}l_{i+1}<l_{i+1}$
+      by 5.3.
+      In this case, for $x\in A_{i+1}$ one has
       $\left|x-\frac{m}{2^{q_{i}-1}}\right|$ $>\frac{2}{3}\left(\frac{1}{2^{k_{i+1}}}\right)-\frac{\alpha l_{i+1}}{2}$
@@
       $>3\alpha l_{i}c$
-    $>\frac{c}{2^{q_{i}-1}}$
-    proving this case and completing the induction.
-Thus, $x_{\infty}\in\operatorname{\widetilde{2-BA}}(c)$ and Alice wins.
-A special case of winning and losing zones given by Theorem 6.4 and Theorem 6.5 can be seen on Figure 5.
-We also consider the special case of $\widetilde{\text{3-BA}}(1/6)$, defined as
-$\widetilde{\text{3-BA}}(1/6):=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{3^{k}}\right|\geq\frac{1}{6\cdot 3^{k}}\text{ for all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\right\}.$
-There, we have some additional bounds:
-Theorem 6.6.
-If
-$\beta\gamma_{A}<\gamma_{B}$
-and
-$\frac{1-\alpha\beta}{\alpha\gamma_{B}}\geq\frac{9\beta}{1-\beta}$
-and
-$\alpha\beta>\frac{1}{7},$
-then the set $\widetilde{\text{3-BA}}(1/6)$ is $(\alpha,\beta)$-losing.
-Proof.
-Let
-$2\varepsilon={\gamma_{B}}-\beta\gamma_{A}.$
-Then, Bob will play in the following way: on his first turn, he plays an interval of length
-$|B_{0}|=\frac{1-\alpha\beta}{\alpha\gamma_{B}}.$
-Then, he will play arbitrarily until some turn $t$, where
-$\frac{1-\alpha\beta}{\alpha\gamma_{B}}\cdot\frac{1}{3^{n+2}}\leq|B_{t}|<(1+\varepsilon)\cdot\frac{1-\alpha\beta}{\alpha\gamma_{B}}\cdot\frac{1}{3^{n+2}}<\frac{1-\alpha\beta}{\alpha\beta\gamma_{A}}\cdot\frac{1}{3^{n+2}}$
-for some $n\in\mathbb{N}$. Then, on turn $t$, he will center his ball at a $c_{t}$ such that
-$c_{t}=\frac{m}{3^{n}}$
-for some $m\in\mathbb{Z}$. On his next move, Bob is always able to play in one of the following two ways:
-  • Bob will center his next interval such that
-    $\left|c_{t+1}-\frac{m}{3^{n}}\right|<\frac{1}{6\cdot 3^{n}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$ (6.3)
-    and Bob will try to minimize $|c_{t+1}-\frac{m}{3^{n}}|$, or
-  • Bob will center his next interval such that
-    $\left|c_{t+1}-\left(\frac{m}{3^{n}}\pm\frac{1}{3^{n+1}}\right)\right|<\frac{1}{6\cdot 3^{n+1}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$ (6.4)
-    and Bob will try to minimize $|c_{t+1}-\frac{m}{3^{n}}\mp\frac{1}{3^{n}+1}|$.
-We show that this is always possible. First, we show that Bob can play his strategy on turn $t$. The length condition is equivalent to
-$0\leq\left\{\log_{3}\frac{|B_{0}|\alpha\gamma_{B}}{1-\alpha\beta}+t\log_{3}\alpha\beta\right\}<\log_{3}(1+\varepsilon)$
-and by the Equidistribution Theorem, there are infinitely many values of $t$ that satisfy the inequality. Then, we have
-$|B_{t}|\geq\frac{1-\alpha\beta}{\alpha\beta\gamma_{A}}\cdot\frac{1}{3^{n+2}}\geq\frac{\beta}{1-\beta}\cdot\frac{1}{3^{n}}$
-so $(1-\beta)|A_{t-1}|\geq 1/3^{n}$ and therefore Bob can always play an interval of the form $m/3^{n}$ on turn $t$.
-Now, we show Bob can play his strategy on turn $t+1$. First, we note that since $\alpha\beta>1/7$, we have
-$\frac{1}{6\alpha\beta}$ $<\frac{7}{6}$
-$\frac{|B_{t}|\gamma_{A}}{2(1-\alpha\beta)}<\frac{1}{2\alpha\beta}\cdot\frac{1}{3^{n+2}}$ $<\frac{7}{6\cdot 3^{n+1}}$
-$\frac{|B_{t}|\gamma_{A}}{2}+\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$ $<\frac{1}{3^{n+1}}+\frac{1}{6\cdot 3^{n+1}}$
-$\frac{m}{3^{n}}+\frac{|B_{t}|\gamma_{A}}{2}$ $<\frac{m}{3^{n}}+\frac{1}{3^{n+1}}+\frac{1}{6\cdot 3^{n+1}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)},$
-which implies that Bob is able to ensure that $c_{t+1}<\frac{m}{3^{n}}+\frac{1}{3^{n+1}}+\frac{1}{6\cdot 3^{n+1}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$.
-By similar logic, Bob can ensure that $c_{t+1}>\frac{m}{3^{n}}-\frac{1}{3^{n+1}}-\frac{1}{6\cdot 3^{n+1}}+\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$.
-Furthermore, we note that
-$\frac{|B_{t}|\alpha\gamma_{B}}{1-\alpha\beta}$ $>\frac{1}{3^{n+2}}$
-$|B_{t}|\alpha\left(1-\beta-\frac{\beta\gamma_{A}}{1-\alpha\beta}\right)$ $>\frac{1}{3^{n+2}}$
-$\frac{1}{3^{n+2}}+\frac{|B_{t}|\alpha\beta\gamma_{A}}{1-\alpha\beta}$ $<(1-\beta)\alpha|B_{t}|$
-$\left|\left(\frac{m}{3^{n}}\pm\left(\frac{1-\frac{1}{6}}{3^{n+1}}+\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}\right)\right)-\left(\frac{m}{3^{n}}\pm\left(\frac{1}{6\cdot 3^{n}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}\right)\right)\right|$ $<(1-\beta)\alpha|B_{t}|,$
-which implies that Bob is always able to satisfy either 6.3 or 6.4.
-Finally, we show this strategy is winning.
-This follows from Lemma 6.3: if 6.3 is true, then in finitely many $k$ turns, $|B_{t+k}|\subset(\frac{m-1/6}{3^{n}},\frac{m+1/6}{3^{n}})$.
-Since Bob can play this way for infinitely many sets of two consecutive turns, Bob wins.
-∎
-Conjecture 6.7.
-If
-$\beta\gamma_{A}\geq\gamma_{B},$
-then the set $\widetilde{\text{3-BA}}(1/6)$ is $(\alpha,\beta)$-winning. If not, then it is $(\alpha,\beta)$-losing.
+      $>\frac{c}{2^{q_{i}-1}},$
+      proving the desired inequality in this last case. This completes the induction, and thus the whole proof.
 Acknowledgements
@@
 Dmitry Kleinbock for the project proposal and his valuable suggestions in research paths.
+We would like to thank Prof.
+Lior Fishman and Prof.
+Stephen Jackson for fruitful discussions of the Schmidt game determinacy problems.
 We would also like to thank Prof.